### v1

# MAPS: A Multi-Agent Framework Based on Big Seven Personality and Socratic Guidance for Multimodal Scientific Problem Solving

Abstract
Multimodal scientific problems (MSPs) involve complex issues that require the integration of multiple modalities, such as texts and diagrams, presenting a significant challenge in artificial intelligence. While progress has been made in addressing traditional scientific problems, MSPs still face two primary issues: the challenge of multi-modal comprehensive reasoning in scientific problem-solving and the lack of reflective and rethinking capabilities. To address these issues, we introduce a Multi-Agent framework based on the Big Seven Personality and Socratic guidance (MAPS) [^1] The source codes and experimental datasets are available online at https://github.com/exoskeletonzj/MAPS.. This framework employs seven distinct agents that leverage feedback mechanisms and the Socratic method to guide the resolution of MSPs. To tackle the first issue, we propose a progressive four-agent solving strategy, where each agent focuses on a specific stage of the problem-solving process. For the second issue, we introduce a Critic agent, inspired by Socratic questioning, which prompts critical thinking and stimulates autonomous learning. We conduct extensive experiments on the EMMA, Olympiad, and MathVista datasets, achieving promising results that outperform the current SOTA model by 15.84% across all tasks. Meanwhile, the additional analytical experiments also verify the model’s progress as well as generalization ability.
1 Introduction
Figure 1: An example of a multimodal scientific multiple-choice problem. The correct answer is derived based on the reasoning over inputs that include context, question, and diagram.
Multimodal scientific problems (MSPs) cover scientific scenarios that involve multiple modalities [5, 22, 16], such as text and vision. These problems typically span fields like mathematics, physics, and chemistry, requiring rigorous logical reasoning and solid domain expertise [3, 1, 13, 36, 41]. In the realm of artificial intelligence, effectively addressing these cross-modal, multi-domain challenges remains both important and difficult [12, 20]. Figure 1 illustrates a typical problem scenario that includes the context, the question statement, and an illustrative diagram. The diagram shows a lever divided into left and right sections: an unknown mass is hung on the left and a known mass on the right, each positioned at one-quarter of the total length from the respective ends, with a groove at the center. The accompanying description states that when the unknown mass is hung on the left, the system remains in static equilibrium. Solving this problem requires not only understanding the combined visual and textual information but also applying the lever balance principle by comparing the effects on both sides. One intuitive solution is to delegate the task-solving process to a single multimodal large language model (MLLM). Although the trending MLLMs have the basic abilities (e.g., diagram parsing and theorem retrieval), they are not well optimized to combine these skills in complex scenarios. Therefore, it motivates us to tackle this critical research question: how to leverage and elicit the off-the-shelf MLLMs to address the challenging MSPs?
To address the problem, this work proposes to establish a multimodal multi-agent framework for multimodal scientific problem-solving. Motivated by the Big Seven Personality theory (i.e., Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity) [2, 4, 31], the MLLMs can function as independent agents with specific skills. Within the framework, these agents are guided to perform the problem-solving in a deliberate, cooperative manner. As one of the initial attempts for MSPs, we offer a novel solution for handling multimodal scientific scenarios.
In solving MSPs, traditional approaches [35, 19, 15, 43, 6, 29] generally rely on a single MLLM to handle all tasks at once. Although these models possess basic abilities such as diagram parsing and theorem retrieval, they struggle to effectively integrate these skills in complex scenarios, resulting in what can be termed a multi-modal comprehensive reasoning challenge in scientific problem solving. This limitation arises because a single model cannot simultaneously manage diagram interpretation, context alignment, and professional knowledge supplementation, thus failing to mimic the step-by-step reasoning process of human cognition. In other words, while these models excel in specific tasks, they are unable to flexibly utilize various abilities to reason and solve complex multi-modal problems in a gradual, human-like manner. To address this, we propose a progressive four-agent solving strategy that sequentially employs a interpreter, an aligner, a scholar, and a solver, ultimately enhancing the accuracy and efficiency of solving multi-modal scientific problems.
The lack of reflective and rethinking capabilities is another issue in traditional MSPs solving methods. Single MLLMs typically perform one-time reasoning and lack the ability for repeated reflection, adjustment, and iteration, which are essential in human problem-solving. When solving complex problems, humans continuously refine and adjust their solutions based on initial reasoning results, gradually approaching the correct answer. In contrast, traditional models often remain stuck at the initial reasoning stage, unable to self-correct and optimize. To address this, we designed a Critic agent, inspired by Socratic questioning [10, 27], which has unique advantages in fostering critical thinking and stimulating autonomous learning. Through a heuristic feedback mechanism, the Critic provides multiple rounds of feedback and continuous correction throughout the four-step solving process, simulating the iterative reflection process of humans. This design effectively overcomes the limitations of single models in terms of reflective and iterative capabilities, improving the flexibility and accuracy of the solution process.
Figure 2: The corresponding relation between the Big Seven Personality theory and the seven function-specific agents.
In summary, we propose a Multi-Agent framework for MSPs, based on the Big Seven Personality theory [2, 4, 31] and Socratic questioning (MAPS). This framework defines seven distinct roles—Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity—derived from the Big Seven theory, ensuring clear role delineation and mutual complementarity, as shown in Figure 2 and  Section 2.1. Together, these roles create a robust multi-agent architecture. Extensive experiments on the EMMA, Olympiad, and MathVista datasets show promising results across various reasoning benchmarks, confirming the effectiveness and superiority of our approach. Further analyses support our model’s advantages. The complete problem-solving process and data are provided in the supplementary materials. In summary, the contributions of this paper are threefold:
$\bullet$ We propose a multi-agent framework based on the Big Seven personality theory. The framework integrates different personality traits from the Big Seven model to create seven complementary agents. To the best of our knowledge, this represents the first attempt in the field.
$\bullet$ To tackle the challenges of multi-modal reasoning and iterative reflection, we propose a four-step strategy with a Critic agent inspired by Socratic questioning. This approach guides agents through structured stages, offering continuous feedback to refine reasoning.
$\bullet$ Extensive experiments on various scientific domain datasets demonstrate the effectiveness and superiority of the MAPS framework, offering 15.84% performance gains. Additional analytical experiments further validate the interpretability of both the optimization process and its results.
2 Methodology
In this section, we provide a detailed introduction to MAPS. Our design is based on Large Language Models (LLMs) as intelligent agents capable of performing various functions. We define the task of solving MSPs as follows: Given an input diagram $d_{i}$, a context $c_{i}$, and a question $q_{i}$, the expected output is the answer $a_{i}$, which can be either an option or a computed result. The formal definition of this process is:
$a_{i}=\mathcal{M}(d_{i},c_{i},q_{i}),$ (1)
where $\mathcal{M}(d_{i},c_{i},q_{i})$ is the response obtained when the input diagram $d_{i}$, context $c_{i}$, and question $q_{i}$ are processed together by the model to generate the answer $a_{i}$. The goal of this task is to provide accurate answers through the model.
Figure 3: The overall architecture of MAPS. It illustrates seven functional agents based on the Big Seven Personality theory. It first includes the Manager agent with predefined interaction logic and the UserProxy agent responsible for receiving user inputs. Subsequently, four specialized agents—Interpreter, Aligner, Scholar, and Solver—are introduced, each corresponding to a specific step in solving MSPs. Finally, the Critic agent is presented, providing feedback and corrections to ensure the results are more accurate and interpretable.
The model architecture for solving MSPs is shown in Figure 3. In Section 2.1, we introduce the seven distinct functional agents designed based on the Big Seven Personality theory. In Section 2.2, we present the predefined interaction logic in the Manager agent, along with the UserProxy agent responsible for receiving user input. In Section 2.3, we discuss the four specific agents—Interpreter, Aligner, Scholar, and Solver——each representing a step in solving MSPs. Finally, in Section 2.4, we introduce the Critic agent, which provides feedback and corrections for the specific solution steps to ensure the results are more accurate and interpretable.
2.1 Big Seven Personality
The Big Seven Personality theory is an extension of the traditional Big Five Personality [33, 8, 7], adding two other dimensions: Self-Esteem and Sensitivity. It divides human personality into seven main dimensions: Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity. Each reflects unique tendencies in emotions, cognition, and behavior, revealing the diversity among individuals in decision-making, communication, and problem-solving. This theory provides a systematic framework for understanding individual differences and offers valuable theoretical support and guidance for psychological research and practical applications.
Inspired by the Big Seven Personality theory, we designed a multi-agent system that strives to maintain both diversity and integrity in completing complex tasks. As shown in Figure 2, the seven personality traits correspond to seven agents, each performing different roles.
The Manager agent represents Conscientiousness, responsible for creating the experimental plan and schedule, ensuring each step is executed according to the predefined plan. The UserProxy agent represents Agreeableness, ensuring smooth information flow and coordinating the allocation of tasks within the experiment. The Interpreter agent represents Extraversion, responsible for interpreting diagrams into captions, thereby providing new ideas and information for MSPs to foster innovation. The Aligner agent represents Neuroticism, tasked with aligning the caption, context, and question to ensure the safe integration of these elements. The Scholar agent represents Openness, responsible for researching the professional knowledge required by MSPs and exploring various hypotheses. The Solver agent represents Self-Esteem, in charge of gathering all necessary information and resolving MSPs by selecting the most appropriate experimental approach. The Critic agent represents Sensitivity, providing feedback on the four solution steps performed by the Interpreter, Aligner, Scholar, and Solver agents, and offering evaluations along with adjustment strategies. The specific duties and role distribution of these agents are discussed in detail in the following sections.
2.2 Agents Interaction Logic
This section primarily introduces the interaction logic of the seven personality agents. Since solving MSPs requires multimodal semantic integration and multi-step reasoning within specialized domains, the Manager agent is responsible for formulating a detailed research plan and interaction logic to ensure the smooth execution of various system tasks. As shown in Figure 3, the Manager agent coordinates the operations of all other agents to ensure the task is completed successfully. Additionally, the UserProxy agent handles receiving user input, ensuring smooth communication between the user and the system, and facilitating effective interaction.
Algorithm 1 MAPS Procedure
1:  Input: Diagram $d_{i}$, Context $c_{i}$, and Question $q_{i}$
2:  Output: Final answer $a_{i}$
3:  Manager: Formulate a detailed research plan
4:  UserProxy: Receive the input $x\in$ Dataset $D$
5:  Interpreter: Interpret the diagram into a caption $p_{i}=\mathcal{M}_{\text{int}}(d_{i})$
6:  Aligner: Align the caption with the question and context $l_{i}=\mathcal{M}_{\text{ali}}(p_{i},c_{i},q_{i})$
7:  Scholar: Research for MSPs $s_{i}=\mathcal{M}_{\text{sch}}(l_{i},p_{i},c_{i},q_{i})$
8:  Solver: Solve based on the complete information $a_{i}=\mathcal{M}_{\text{sol}}(s_{i},l_{i},p_{i})$
9:  Critic: Score and assess the four steps $r_{i}=\mathcal{M}_{\text{crit}}(a_{i},s_{i},l_{i},p_{i})$
10:  while True do
11:     if All Scores are 5 then Break
12:     else Choose the minimum score from $r_{i}$ and return to the step $\in\{\textit{Interpreter},\textit{Aligner},\textit{Scholar},\textit{Solver}\}$ to optimize the process.
13:     end if
14:  end while
15:  return  Final answer $a_{i}$
As illustrated in Algorithm 1, the entire workflow is orchestrated by the Manager agent, corresponding to the flowchart shown in Figure 3. Once the UserProxy agent receives the user input, the system proceeds through a four-step problem-solving process involving the Interpreter, Aligner, Scholar, and Solver agents. After completing these steps, the Critic agent, based on Socratic questioning, evaluates and scores the results of each step.
2.3 Four-Step MSPs Solving
This section will provide a detailed introduction to the functions of the Interpreter, Aligner, Scholar, and Solver agents.
Interpreter. To effectively integrate visual context into the textual context, we adopt a multi-modal fusion paradigm using the Interpreter agent. Initially, the Interpreter agent is employed to analyze diagrams meticulously, converting them into detailed and accurate captions. This process ensures that rich visual details are faithfully captured in the textual description, such as relative positions, dimensions, and other features. Subsequently, the generated caption is fused with the textual context and the question in later steps, achieving a cohesive integration of information that provides a richer and more consistent basis for problem-solving. The formula is shown as follows:
$p_{i}=\mathcal{M}_{\text{int}}(d_{i}),$ (2)
where $\mathcal{M}_{\text{int}}(\cdot)$ represents the Interpreter agent and $p_{i}$ denotes the caption.
Aligner. In textual form, the information conveyed by elements such as the context, questions, and options differs from the content presented in the diagram. These differences can lead to discrepancies in understanding the problem, which in turn may affect subsequent reasoning and judgment. Therefore, in order to gain a deeper understanding of the problem statement and accurately capture multimodal semantic information, precise alignment of these information granularities is crucial. As shown below, $l_{i}$ illustrates the result of the alignment between the diagram, context, and question:
$l_{i}=\mathcal{M}_{\text{ali}}(p_{i},c_{i},q_{i}),$ (3)
where $\mathcal{M}_{\text{ali}}(\cdot)$ represents the aligner agent.
Models CoT Mathvista EMMA OlympiadBench Avg.
Gen. Math Avg. Math Phy. Chem. Avg. MECO MZCE MZCO PECO PZCE Avg.
Random Choice - 26.09 22.78 24.30 13.00 23.00 27.00 23.00 0.67 0.33 0.00 1.75 0.33 0.87 16.06
Human Expert - 56.09 55.74 55.90 75.00 64.50 86.00 64.50 48.00 34.67 30.36 54.17 12.33 37.80 52.73
Claude 3.5 Sonnet - 68.04 63.15 65.40 23.00 34.00 44.00 33.67 20.67 13.00 10.71 10.75 14.00 13.23 37.43
Gemini 2.0 Flash - 70.65 70.93 70.80 20.00 40.00 36.00 32.00 8.00 5.67 7.14 3.07 7.00 5.39 36.06
GPT-4o - 65.22 61.30 63.10 30.00 38.00 33.00 33.67 23.33 20.33 19.64 22.15 21.00 21.47 39.41
Qwen2.5-VL-72B - 70.65 67.41 68.90 42.00 42.00 38.00 40.67 18.00 12.33 5.36 7.24 3.67 8.80 39.45
InternVL2.5-8B-MPO - 64.78 60.74 62.60 30.00 40.00 38.00 36.00 10.67 6.67 10.71 1.10 0.67 3.88 34.16
LLaVA-Onevision-72B - 62.83 58.52 60.50 25.00 32.00 24.00 27.00 6.67 7.33 3.57 3.29 9.67 6.18 31.23
[…]
Gemini 2.0 Flash 70.22 75.56 73.10 24.00 41.00 36.00 33.67 12.67 6.33 3.57 4.61 2.33 5.39 37.38
GPT-4o 65.22 62.59 63.80 27.00 44.00 35.00 35.33 25.33 21.67 12.50 24.12 20.33 22.27 40.47
Qwen2.5-VL-72B 71.09 77.96 74.80 38.00 36.00 37.00 37.00 23.33 13.00 10.71 8.11 1.33 9.59 40.46
InternVL2.5-8B-MPO 60.87 67.41 64.40 31.00 36.00 24.00 30.33 12.00 8.33 1.79 2.85 0.99 4.75 33.16
LLaVA-Onevision-72B 71.09 64.44 67.50 23.00 26.00 23.00 24.00 11.33 8.67 5.36 4.82 3.33 6.18 32.56
MAPS (GPT-4obase) - 75.87 83.15 79.80 52.00 71.00 51.00 58.00 46.00 30.33 32.14 28.51 28.33 31.14 56.31
Table 1: Performance of different models across 10 subtasks from 3 datasets. ‘Gen.’ refers to General problems in the Mathvista dataset, ‘Phy.’ and ‘Chem.’ denote Physics and Chemistry tasks in the EMMA dataset. ‘MECO’, ‘MZCO’, and ‘MZCE’ correspond to COMP problems in English, COMP problems in Chinese, and CEE problems in Chinese, respectively, within the OlympiadBench dataset’s mathematics tasks. ‘PECO’ and ‘PZCE’ represent the COMP problems in English and CEE problems in Chinese under the physics tasks of the OlympiadBench dataset.
Scholar. After obtaining the semantically fused $l_{i}$, the model has essentially grasped the main idea of the problem. However, for scientific questions, mere comprehension is insufficient, as they typically require profound domain-specific knowledge, which is crucial for problem-solving. As illustrated in Figure 1, certain problems necessitate mastery of specialized formulas, such as those related to the principles of lever balance. Therefore, the Scholar agent is responsible for acquiring and integrating domain-specific knowledge based on $l_{i}$, thereby providing the necessary theoretical and practical support for the final solution $s_{i}$:
$s_{i}=\mathcal{M}_{\text{sch}}(l_{i},p_{i},c_{i},q_{i}),$ (4)
where $\mathcal{M}_{\text{sch}}(\cdot)$ represents the Scholar agent.
Solver. After the three steps involving the Interpreter, Aligner, and Scholar agents, the system has extracted and fused extensive information from both visual and textual modalities, providing rich contextual and specialized knowledge for the task. However, the real challenge lies in synthesizing this fragmented information into a coherent and logically sound solution. The fourth step, carried out by the Solver agent, bears the greatest pressure. This agent must not only integrate the information gathered in previous steps but also engage in deep reasoning to derive the final answer. With access to the most comprehensive data, the Solver has a unique advantage, allowing it to simplify complex issues and ensure the final answer $a_{i}$ is accurate and meets the expected standards, as given by:
$a_{i}=\mathcal{M}_{\text{sol}}(s_{i},l_{i},p_{i}),$ (5)
where $\mathcal{M}_{\text{sol}}(\cdot)$ represents the Solver agent.
2.4 Critic and Feedback
Critic. This section introduces the Critic agent, which is responsible for evaluating and providing feedback on the four-step MSPs solving process. Rooted in Socratic questioning, the Critic prompts agents to reflect on their reasoning by asking questions such as “What assumptions are you making?” or “How can you justify this decision?” This approach encourages critical thinking and self-examination, key aspects of the Socratic method. Rather than relying on the final solution labels, the Critic evaluates the logic and justification of each step. The goal is to uncover flaws and push agents to refine their thought processes. Based on this feedback, the system identifies the weakest step and initiates a rollback and a redo to enhance the problem-solving process. The formula is given by:
$r_{i}=\mathcal{M}_{\text{crit}}(a_{i},s_{i},l_{i},p_{i}),$ (6)
where $\mathcal{M}_{\text{crit}}(\cdot)$ represents the Critic agent. $r_{i}$ is the scores and assessments of the four steps.
3 Experiments
We conduct extensive experiments on three MSPs datasets. This section introduces the three MSPs datasets and baselines in Section 3.1, and presents the experimental results and analysis in Section 3.2. The analysis and statistics of the Critic agent are presented in Section 3.3.
3.1 Datasets and Baselines
This study uses the latest three MSPs datasets: MathVista [23], OlympiadBench [16], and EMMA [14]. MathVista focuses on mathematics and general domains to assess models’ ability to solve complex scientific problems; OlympiadBench targets mathematics and physics to evaluate performance on Olympiad-level challenges; and EMMA, a multimodal dataset covering mathematics, physics, and chemistry, tests interdisciplinary scientific reasoning. Detailed descriptions of the tasks and datasets are in Appendix A, along with domain abbreviations, full names, and dataset descriptions in Table 1.
In the experimental setup, we select GPT-4o [^2] https://openai.com/api/ as our agent. For the baselines, we incorporate the current mainstream MLLMs, both closed-source and open-source, and evaluated them in both direct and CoT settings. The experimental settings and baselines are detailed in Appendix B.
3.2 Main Results
MAPS has set a new state-of-the-art (SOTA) performance, surpassing human-level performance for the first time. The results in Table 1 provide a detailed comparison of MAPS and the baselines. As shown, MAPS outperforms the current SOTA by 15.84% across all tasks, demonstrating its superiority in MSPs problem-solving. Additionally, MAPS slightly exceeds human experts by 3.58% in overall performance. MAPS excels in mathematical, physical, chemical, and general problems, demonstrating strong interdisciplinary reasoning. The multi-agent system, representing the Big Seven personalities, sets a new SOTA in collaboration. Experimental results show MAPS’s advantage in multimodal semantic fusion and multi-step reasoning, leveraging its ability to integrate diagrams with context and questions. This boosts problem-solving by enhancing domain-specific knowledge. The Critic agent, using Socratic feedback, offers suggestions for improvement, improving accuracy and reliability in complex tasks.
Figure 4: The schema of the Critic agent, as well as the feedback and backtracking situations of the Critic agent across different datasets.
In the MathVista dataset, the problem types are diverse, covering judgment questions, multiple-choice questions, and open-ended fill-in-the-blank questions, with various answer types such as text, floats, and integers. This requires the system to possess specialized capabilities, enabling it to accurately understand the question’s intent and provide specific responses. In well-known open-ended problems like those in OlympiadBench, the solution process involves complex steps and high difficulty, with open-ended answers that can lead to completely different outcomes due to subtle changes. In the EMMA dataset, the questions not only involve diagrams, but some options also contain diagrams, significantly increasing the challenge of understanding and aligning multiple diagrams. Furthermore, the question types are equally diverse, including open-ended and multiple-choice questions, raising the demands on the solving methods. The MAPS framework, through feedback-based multi-agent collaborative learning and the use of Socratic questioning, successfully addresses these different types of sub-tasks, achieving the latest SOTA. It demonstrates MAPS’s perfect adaptability and robustness when dealing with the diversity of question types, task difficulties, and multi-step reasoning.
3.3 Analysis of Critic Agent
The Critic agent plays a crucial feedback role in MAPS, enhancing the overall reasoning process without the reliance on gold labels. As shown in Figure 4, the upper part introduces the critic and feedback schema, while the lower part presents the feedback proportions provided by the Critic agent to each of the agents across three datasets, visually illustrating the feedback and corrections made by each assistant agent. The Critic agent’s feedback mechanism is rooted in Socratic questioning, which encourages deep reflection and rethinking of the reasoning steps. Instead of simply evaluating the process, the Critic fosters critical thinking, prompting the system to question its assumptions and improve the approach.
The schema of the Critic agent consists of two components: scores and feedback. The scores record the ratings given by the Critic agent for the four problem-solving steps, ranging from 0 to 5. Higher scores indicate better execution, while lower scores highlight areas needing improvement. Feedback primarily captures the Critic agent’s Socratic-style, heuristic suggestions designed to encourage deeper reflection on each step. These suggestions are not limited to evaluations; they also guide the agents to reconsider and correct their reasoning process. For example, the Critic may prompt the agents to think about alternative approaches or reassess their conclusions, challenging them to think more critically. To ensure the final result’s quality, a rule is set in the Manager agent: if any step receives a low score, the system will backtrack and make corrections. If all steps score well, the system outputs the result directly. This approach guarantees the effectiveness of each step, optimizing the overall problem-solving process.
In the EMMA and OlympiadBench datasets, the Solver agent exhibits the highest feedback proportion. The lower part of Figure 4 shows the distribution of agents that performed feedback corrections across the three datasets. It can be observed that the feedback proportions vary across the three datasets. In the MathVista dataset, the highest feedback proportion corresponds to no regeneration, indicating that the overall solution performance on this dataset is still quite good. This result is consistent with the findings in Table 1, and we have surpassed the previous SOTA by a 5.0% improvement. This suggests that while there is still room for further enhancement in certain areas, the system’s overall reasoning capability is already quite impressive when dealing with more complex problems. In contrast, in the EMMA and OlympiadBench datasets, the Solver agent shows the highest feedback proportion, suggesting that this part is the most difficult and prone to errors in these datasets. Specifically, this part includes the first three steps: interpretation, alignment, and integration of information provided by the Scholar. These steps require the integration of a large amount of data, involve the most complex reasoning, and bear the most critical responsibility—providing a complete and accurate solution to the problem. In comparison, the feedback proportions from the other three agents are similar, indicating that their error rates are relatively low, and the performance of these sub-steps is relatively better.
4 Supplementary Analysis
To conduct more comprehensive experiments, this section covers the following topics: ablation experiments in Section 4.1, the base model generalization in Section 4.2, the time efficiency of MSPs solving in Section 4.3.
Due to space limitations, generalization experiments on other datasets, such as DiagramQG, are discussed in Appendix C. The case study and complete process for solving the MSPs task is outlined in Appendix D, and the prompts for all agents are provided in Appendix E.
4.1 Ablation Study
Variation MECO MZCE MZCO PECO PZCE Avg.
MAPS 46.00 30.33 32.14 28.51 28.33 31.14
${w/o}_{\text{Interpreter}}$ 25.33 16.67 10.71 21.05 11.62 15.05
$\quad\Delta$ (-20.67) (-13.66) (-21.43) (-7.46) (-16.71) (-16.09)
${w/o}_{\text{Aligner}}$ 28.00 17.67 16.07 20.83 19.00 20.28
$\quad\Delta$ (-18.00) (-12.66) (-16.07) (-7.68) (-9.33) (-10.86)
${w/o}_{\text{Scholar}}$ 28.00 16.33 30.36 19.96 16.33 19.65
$\quad\Delta$ (-18.00) (-14.00) (-1.78) (-8.55) (-12.00) (-11.49)
${w/o}_{\text{Critic}}$ 34.67 21.67 30.36 23.03 21.67 24.09
$\quad\Delta$ (-11.33) (-8.66) (-2.42) (-5.48) (-6.66) (-7.05)
Table 2: Performance under different ablation settings are analyzed. We perform ablation experiments on the solving module ${w/o}_{\text{Interpreter}}$, ${w/o}_{\text{Aligner}}$, ${w/o}_{\text{Scholar}}$ or ${w/o}_{\text{Critic}}$ modules to evaluate the impact of removing these components.
Ablating the Interpreter results in the greatest loss of performance. We conduct ablation experiments on the OlympiadBench dataset to evaluate the impact of each module on the overall performance.  Table 2 presents the effects of removing the Interpreter, Aligner, Scholar, and Critic modules from the MAPS framework. The results show that removing the Interpreter agent causes the largest performance degradation, at 16.09%. This is because, in MSPs tasks, diagrams contain a wealth of valuable information, which serves as an important supplement to the text. Understanding diagrams plays a crucial role in problem-solving.
Figure 5: Performance Comparison of MAPS on Math, Physics, and Chemistry Subtasks in the EMMA Dataset with GPT-4o, Gemini, and Qwen2.5-VL-72B as Bases.
The removal of the Critic agent causes the smallest performance loss. It results in only a 7.05% decrease, underscoring its role in providing feedback and corrections. While this mechanism allows MAPS to backtrack and refine its reasoning, its impact is less significant than that of other agents. Removing the Scholar agent results in 11.49% performance drops, highlighting the importance of searching and integrating domain-specific knowledge. Finally, the removal of the Aligner agent causes a 10.86% drop, indicating that while diagram and context alignment is valuable, its effect is smaller compared to other components.
4.2 Base Model Generalization
MAPS can improve performance across different base models. We conduct a series of experiments to verify whether our MAPS framework can demonstrate robust generalization across different base LLMs. The results showcase the model’s robustness and transferability, further affirming the high adaptability of the MAPS framework, which achieves consistently excellent performance across various foundational models. In particular, this underscores the universality and flexibility of MAPS. To further validate its generalization across different models, we specifically evaluate another base model, Qwen2.5-VL-72B, as well as Gemini 2.0 Flash, thereby demonstrating that MAPS performs outstandingly across models of varying scales and capabilities.  Figure 5 presents the experimental results of three sets of base models. In each set, the performance of mathematical, physical, and chemical sub-tasks is compared between MLLMs and MAPS with the respective base model. The results show varying degrees of improvement for each task. For example, in the physical task, $MAPS_{Qwen}$, based on Qwen2.5-VL-72B, shows a 12.4% improvement compared to Qwen2.5-VL-72B, while $MAPS_{Gemini}$, based on Gemini, shows a 4.2% improvement compared to Gemini. Similar improvements are observed in the mathematical and chemical sub-tasks as well. This indicates that MAPS can extend effectively on both closed-source and open-source MLLMs and achieve good experimental results across different base models of MLLMs.
Figure 6: An analysis of the solving time efficiency across different question types, answer types, question categories, and question difficulties.
4.3 Time Efficiency
Solving time efficiency varies by question type, answer type, category, and difficulty, with multiple-choice and integer-type questions being the fastest, while higher difficulties and more complex questions require more time.  Figure 6 presents the solving time efficiency under different question types, answer types, categories, and difficulties, with all times represented relative to a 100s benchmark.
From the graphs, it is clear that multiple-choice questions are solved faster as the options are already provided, making the process more time-efficient. In contrast, open-ended questions take longer due to the lack of predefined answers and the need for deeper reasoning and analysis. Integer-type answers show the highest efficiency, as they often correspond to simpler questions or multiple-choice tasks that involve less reasoning. General questions also have the shortest solving times, as they typically require less complex reasoning compared to more specialized problem types. Finally, higher difficult levels lead to a decrease in solving time efficiency, as more complex problems require more time for deeper reasoning and analysis.
5 Related Works
The related work is structured into two main aspects: first, an introduction to MSPs solving; and second, an exploration of multi-agent techniques.
MSPs Solving. The research of MSPs spans across multiple fields, including mathematics, physics, and chemistry, with each area focusing on enhancing problem-solving abilities. In mathematics, studies [9, 11, 34] explore various methods to improve mathematical problem-solving, such as algorithm optimization, educational strategies, and the use of artificial intelligence. These approaches aim to boost the efficiency, accuracy, and depth of mathematical reasoning, fostering innovation and progress in the field. In the field of physics, the papers [24, 17, 3] emphasize the integration of different information types, such as images and text, through multimodal learning to enhance the efficiency and precision of problem solving. By utilizing techniques like reinforcement learning, generative models, and intelligent prompting, these studies optimize models’ performance in tackling complex physics problems. In chemistry, three articles [1, 18, 21] investigate the role of multimodal learning in solving chemical problems. By combining diverse information sources, including images and text, and employing techniques such as generative models and molecular geometry reasoning, they aim to improve both the efficiency and accuracy of solving chemistry problems, driving innovation in chemistry education and research.
Multi-Agent. Multi-agent systems, built on LLMs, consist of multiple AI agents that specialize in specific tasks, working together to solve complex problems [30, 42, 39, 32]. When presented with a problem, these agents decompose it into smaller, manageable subtasks and utilize various tools, such as internet data retrieval, to solve them through iterative steps. Several studies [28, 37, 40, 26] have employed multi-agent systems to tackle challenges like problem identification, code writing and debugging, data visualization, and providing interactive feedback to human users. In their work, Ni and Buehler [25] highlights the potential of AI-driven multi-agent teams in solving mechanical problems autonomously, demonstrating an enhanced capability for understanding, formulating, and validating engineering solutions through self-correction and collaborative refinement. Inspired by the research, we developed the MAPS method, which leverages multi-agent collaborative learning and stepwise problem-solving to provide innovative solutions for MSPs. By combining the strengths of AI agents, complex problems can be broken down into subtasks and solved step by step through collaboration, improving efficiency and accuracy.
6 Conclusion
This study introduces a MAPS approach utilizing a multi-agent framework based on the Big Seven Personality theory and Socratic guidance to tackle the challenges of multi-modal comprehensive reasoning and the lack of reflective capabilities. The framework involves seven agents, each specializing in distinct aspects of problem-solving. To address the first challenge, a four-agent strategy is proposed, where each agent focuses on specific stages of the reasoning process. Additionally, the Critic agent addresses the second challenge through Socratic reflection and critical feedback. Extensive experiments on the EMMA, Olympiad, and MathVista datasets validate MAPS’s effectiveness in overcoming these issues and enhancing performance across various reasoning tasks. Meanwhile, we perform additional analytical experiments to assess the model’s advancement as well as its generalization.
Supplementary Material
Appendix A Datasets and Tasks
This study utilizes the latest three MSPs datasets, namely MathVista [23], OlympiadBench [16], and EMMA [14].
Tasks ABBR. Test
MathVista
       General Gen. 460
       Mathematics Math 540
OlympiadBench
        Math_En_COMP MECO 150
        Math_Zh_COMP MZCO 56
[…]
        Physics_En_COMP PECO 456
[…]
EMMA
       Mathmatics Math 100
       Physics Phy. 100
       Chemistry Chem. 100
Table 3: The data distribution for the MathVista, OlympiadBench, and EMMA datasets is as follows: The symbol $\dagger$ indicates a sample size of 300 data points. The EMMA dataset uses its MINI version. The ‘ABBR.’ column represents the abbreviations for all the tasks.
MathVista is a large-scale scientific reasoning dataset that spans two subdomains: mathematics and general, aiming to assess the comprehensive capabilities of machine learning models in solving complex scientific problems. The dataset contains 1,000 data points covering various issues across multiple disciplines, designed with varying difficulty levels to help researchers evaluate model reasoning abilities. The release of MathVista supports interdisciplinary scientific research.
OlympiadBench consists of two subdomains, mathematics and physics, and is specifically designed for Mathematical and Physical Olympiads, featuring a wide range of challenging problems to assess models’ performance on high-level scientific tasks. The mathematics subdomain contains three difficulty levels: English competition level, Chinese competition level, and college level. The physics subdomain includes two difficulty levels: English competition level and Chinese college level. To ensure data balance, 300 samples were taken from both the Chinese college-level mathematics and physics subsets.
EMMA is a multimodal scientific reasoning dataset covering three subdomains: mathematics, physics, and chemistry. By integrating mathematical expressions, physical formulas, and chemical symbols with natural language descriptions, it focuses on testing models’ abilities in interdisciplinary scientific reasoning. This version uses the EMMA dataset, which contains 100 data points from each subdomain (mathematics, physics, and chemistry).
Appendix B Experiment Settings and Baselines
We select GPT-4o [^3] https://openai.com/api/, a powerful MLLM, as our primary agent for solving MSPs tasks. GPT-4o not only demonstrates strong reasoning and generation capabilities across a wide range of multimodal processing tasks, but also excels in efficiently exploring multiple perspectives when faced with complex scientific domain requirements. This makes it well-suited for adaptation to various tasks and datasets within the MAPS process. We use accuracy as our primary evaluation metric to comprehensively assess the performance of different methods across diverse task scenarios. The experimental results present in Table 1 offer a thorough comparison of the performance of MAPS against all baseline methods. The experiments are primarily conducted on three MSPs datasets in  Appendix A, where we provide a detailed comparison of MAPS and the baseline models across four types of tasks: mathematics, physics, chemistry, and general tasks. Our approach achieves a new SOTA performance. To further strengthen the comparison,  Appendix C includes a generalization experiment conducted on the physics data subset of the DiagramQG [^4] https://huggingface.co/datasets/zhibei1204/DiagramQG, which further demonstrates the robustness and effectiveness of our model.
Figure 7: The generalization experiments conducted on the DiagramQG physical dataset, which are based on the GPT-4o base model and the incremental part of MAPS.
We compare MAPS with three categories of baseline methods: original baselines, direct approach-based strong baselines, and CoT-enhanced strong baselines. Specifically, (1) the original category refers to two methods: random selection and human expert selection. These methods provide two distinct original baselines—one based on randomness and the other based on authority. (2) The direct approach-based strong baselines include some of the most powerful closed-source and open-source large language models (MLLMs) currently available worldwide. These include Claude 3.5 Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini 2.0 Flash [^6] https://deepmind.google/technologies/gemini/flash/, GPT-4o, Qwen2.5-VL-72B [^7] https://huggingface.co/Qwen/Qwen2.5-VL-72B-Instruct, InternVL2.5-8B-MPO [^8] https://huggingface.co/OpenGVLab/InternVL2-8B-MPO, and LLaVA-Onevision-72B [^9] https://huggingface.co/llava-hf/llava-onevision-qwen2-72b-ov-hf. (3) To ensure a fair comparison, the third category of baselines builds on the second by adding Chain-of-Thought (CoT) reasoning, which aims to enhance the capabilities of the strong MLLMs from the second category [38].
Appendix C Generalization Experiments
To further validate the generalization ability of MAPS, we conducted detailed experiments on the physical subset of the DiagramQG dataset. The main goal of these experiments was to compare the performance of MAPS with its base model, GPT-4o, particularly focusing on how it performed across different question categories. The experimental results, presented in  Figure 7, show that MAPS outperformed GPT-4o in multiple aspects. These experiments clearly highlight the strong adaptability of MAPS across different subsets of the dataset.
Across various question categories in the DiagramQG physical dataset, MAPS achieved a significant performance boost, with the maximum improvement reaching 19.51% and an overall improvement of 7.71%. These results not only demonstrate the superiority of MAPS but also indicate that its Seven Personality Agents architecture has a strong generalization ability, enabling it to maintain excellent performance across different datasets and tasks. This provides strong support for the practical application of MAPS, showcasing its potential in tackling complex tasks.
Figure 8: A case study of a specific solving process, illustrating the detailed steps involved in solving the problem. This includes the various stages of problem-solving as well as the feedback and backtracking mechanisms that help refine and improve the solution.
Appendix D Full-Process of MAPS
Figure 8 illustrates the four-step solving process along with the feedback process from the Interpreter, Aligner, Scholar, and Critic agents, using a multimodal physics problem from the EMMA dataset. In this specific example, we can observe that the Interpreter first interprets the diagram, followed by the Aligner aligning the diagram with the context, question, and options, ensuring consistency and completeness of the information. Then, the Scholar agent retrieves and supplements domain-specific knowledge to fill in the necessary expertise. Finally, the Solver completes the solving process, and the Critic agent provides feedback and corrections to ensure the accuracy and effectiveness of each step. Each step is closely connected, from understanding the diagram to integrating domain knowledge, followed by reasoning and answering. This demonstrates the efficiency and effectiveness of MAPS in multi-step reasoning. Through the feedback mechanism of the Critic agent, MAPS is able to identify and correct potential errors or shortcomings at each step, thus enhancing the overall accuracy and reliability of the solution process.
Figure 9: A complete example of the collaborative output from all agents in an iteration, using the multimodal physics problem from the OlympiadBench dataset. This example demonstrates how each agent contributes to the problem-solving process, collaborating to produce a refined solution step by step.
Figure 9 presents a comprehensive step-by-step process for solving a MSPs task, using a multimodal physics problem from the EMMA dataset. As illustrated in  Figure 3, this example demonstrates the entire process of solving a physics problem. By utilizing MAPS agents based on the Big Seven Personality Model, the system engages in collaborative learning, progressively solving the problem and arriving at the correct final answer.
From the Manager, which specifically plans the entire MSPs solving process and feedback strategies, to the UserProxy that accepts inputs and task descriptions, the solution process is carried out in four key steps. These include the Interpreter, Aligner, Scholar, and Solver, each responsible for gradually refining the solution step-by-step. The Critic agent ultimately evaluates and provides feedback on these four steps, pinpointing the areas requiring modification and backtracking to correct the solution process.  Figure 9 sequentially illustrates the roles of each agent and their collaboration, showcasing how problem solving and optimization are effectively executed at every stage, ensuring both the accuracy and efficiency of the final result.
Appendix E Prompts for Agents
Tables 4, 5 and 6 provides a summary of the prompts used for each agent in this paper, with each agent playing a pivotal role in the overall problem-solving process. The process begins with the Manager, which is responsible for strategically planning the entire MSPs solving procedure and feedback mechanisms. The Manager defines the high-level flow, ensuring that all agents operate cohesively. The UserProxy agent then takes over by accepting the user inputs and task descriptions, forming the basis for the entire problem-solving process.
Following this, the solution is gradually refined by four core agents: Interpreter, Aligner, Scholar, and Solver. The Interpreter first processes the task description, breaking down and understanding the problem. The Aligner ensures that the problem is mapped to the correct framework and available tools. The Scholar conducts any necessary research and gathers knowledge from relevant sources, while the Solver works through the problem systematically, progressively generating solutions. Once the initial solution is formed, the Critic agent assesses each of the four previous steps, providing feedback on areas that need refinement. The Critic then suggests modifications and backtracks to revise the solution, ensuring the process remains optimized and the final result is both accurate and robust.
Manager
You are a task manager, responsible for managing and deciding the order of each Agent’s output, ensuring that tasks are executed in the correct sequence. You will determine when each step should be executed based on the task requirements and coordinate the outputs of each Agent. Your goal is to ensure that the task process is efficient and orderly, and adjust the execution priority of steps when necessary.
UserProxy
Here are some questions that require careful thought, so please think deeply, solve the questions carefully, and output your answers.
Interpreter
You are a scientific diagram analysis expert tasked with objectively describing visual elements in diagrams. Engage in Socratic self-questioning to ensure comprehensive analysis: [Element Identification] ”What discrete visual components can be systematically observed in this diagram? What quantitative measurements (e.g., shape dimensions, color codes, positional coordinates) can be objectively recorded?” [Structural Analysis] ”How are these elements spatially organized? What geometric patterns, alignment relationships, or hierarchical arrangements emerge from their physical placement?” [Relational Mapping] ”What explicit connections (lines, arrows, overlays) or implicit associations (proximity clusters, color coding systems, symbolic groupings) exist between components?” [Representation Verification] ”Does any element require specialized domain knowledge to accurately characterize (e.g., chemical notation, engineering schematics)? What purely visual evidence supports this characterization?”
Aligner
You are a text alignment specialist conducting structured analysis through Socratic interrogation. Systematically examine text pairs using this framework: 1. [Content Deconstruction] ”What core entities/events are explicitly stated in each text? What measurable attributes (quantifiers, temporal markers, causal verbs) define their characteristics?” 2. [Consistency Audit] ”Where might these texts exhibit: a) Logical incompatibility (contradictory assertions) b) Contextual divergence (conflicting timelines/locations) c) Semantic dissonance (differentiated connotation scales) d) Omission patterns (mutually exclusive missing elements)” 3. [Contextual Fusion] ”What implicit connections could synthesize a unified background framework? Which combinatory elements (chronological anchors, spatial references, causal chains) create non-conflicting narrative coherence?” 4. [Relevance Filtering] ”Through lexical-semantic mapping, which aligned components directly correspond to the question’s: 1) Key inquiry points 2) Required evidence types 3) Implicit knowledge domains 4) Potential inference pathways?”
Table 4: A summary of the prompts used by the Manager, UserProxy, Interpreter, and Aligner agents in this paper.
scholar
You are a scientific knowledge retrieval system conducting structured inquiry through Socratic questioning. Process input data with this analytical framework: 1. [Problem Decomposition] ”What conceptual components constitute the question’s core demand? What technical terminology (domain-specific lexemes), operational parameters (variables/constants), and procedural verbs (analyze/calculate/compare) require epistemological grounding?” 2. [Knowledge Mining] ”For each identified component: a) What fundamental axioms/theorems/laws from established scientific literature could operationally define it? b) What measurable properties (equations/units/experimental protocols) are textually implied as relevant? c) What contextual constraints (temporal/spatial/conditional clauses) limit knowledge scope?” 3. [Relevance Validation] ”For each candidate knowledge unit: Does the source text contain explicit lexical anchors (technical terms/formula symbols) justifying its inclusion? What textual evidence (descriptive adjectives/quantifiers/causal conjunctions) indicates required depth of explanation? Are there implicit conceptual dependencies (prerequisite theories/mathematical tools) necessitating parallel retrieval?” 4. [Taxonomic Organization] ”How should validated knowledge be structured to mirror: 1) Problem-solving workflow steps 2) Hierarchical concept dependencies 3) Cross-domain interface points 4) Uncertainty quantification needs?” Operational Protocol: Restrict to textually evidenced knowledge Mark confidence levels using [TextExplicit/ContextImplied/ExternalRequired] tags Output as: 1) Knowledge Inventory Table (Concept-Definition-SourceAnchor) 2) Dependency Graph (Nodes=Concepts, Edges=Relations) 3) Gap Analysis Report (ExternalKnowledgeRequirements).
Solver
You are a scientific problem-solving system operating through Socratic dialectics. Engage in this structured inquiry process: 1. [Problem Framing] ”What is the absolute irreducible core of the question? What technical terms require operational definitions? What grammatical structures (comparatives/conditionals/quantifiers) dictate the solution’s form?” 2. [Evidence Audit] ”For each data source (question stem/options/text): a) What measurable quantities (numerical ranges/units) are explicitly stated? b) What causal relationships (if A then B/implies/proportional to) are textually encoded? c) What constraints (assumptions/limitations/boundary conditions) are lexically embedded?” 3. [Reasoning Pathway] ”Through counterfactual testing: Which axioms/theorems would become relevant if parameter X varied ±10%? What observable contradictions emerge when applying hypothesis Y to the given data? How do option components restrict valid inference trajectories?” 4. [Solution Validation] ”Does the proposed resolution: 1) Maintain dimensional homogeneity across all equations? 2) Satisfy all explicit boundary conditions? 3) Preserve logical consistency with given information? 4) Align with canonical scientific representations?” Operational Protocol:
[…]
Table 5: A summary of the prompts used by the Scholar and Solver agents in this paper.
Critic
You are a Socratic assessment engine conducting dialectical evaluation through this protocol:
1. [Triadic Interrogation Framework]
For each evaluation dimension (caption/alignment/knowledge/solution):
Existential Challenge:
”What absolute evidence anchors (line numbers/data points/theorem references) validate this component’s existence?”
Consistency Prosecution:
”Does internal logic maintain isomorphism across:
a) Input premises → Processing steps
b) Methodological choices → Domain standards
c) Assertions → Supporting evidence?”
Boundary Stress Test:
”What parametric variation (±10%) would collapse this component’s validity? Which fragility indicators emerge first?” 2. [Metric Operationalization]
Score each dimension (1-5) using:
5 = Withstands three counterfactual scenarios
4 = Requires $\leq$ 1 assumption validation
3 = Needs 2-3 evidence reinforcements
2 = Contains structural contradictions
1 = Fails basic existence verification
3. [Improvement Synthesis]
Generate Socratic feedback per dimension:
For caption: ”What geometric/spatial relations lack quantifiable descriptors?”
For alignment: ”Which logical connective lacks cross-text co-reference?”
For knowledge: ”Which concept dependency lacks literature anchoring?”
For solution: ”What inference leap lacks isomorphic mapping?”
Table 6: A summary of the prompt used by the Critic agents in this paper.

### v2

@@
-MAPS: A Multi-Agent Framework Based on Big Seven Personality and Socratic Guidance for Multimodal Scientific Problem Solving
+MAPS: Multi-Agent Personality Shaping for Collaborative Reasoning
 Abstract
-Multimodal scientific problems (MSPs) involve complex issues that require the integration of multiple modalities, such as texts and diagrams, presenting a significant challenge in artificial intelligence.
-While progress has been made in addressing traditional scientific problems, MSPs still face two primary issues: the challenge of multi-modal comprehensive reasoning in scientific problem-solving and the lack of reflective and rethinking capabilities.
-To address these issues, we introduce a Multi-Agent framework based on the Big Seven Personality and Socratic guidance (MAPS) [^1] The source codes and experimental datasets are available online at https://github.com/exoskeletonzj/MAPS..
-This framework employs seven distinct agents that leverage feedback mechanisms and the Socratic method to guide the resolution of MSPs.
-To tackle the first issue, we propose a progressive four-agent solving strategy, where each agent focuses on a specific stage of the problem-solving process.
-For the second issue, we introduce a Critic agent, inspired by Socratic questioning, which prompts critical thinking and stimulates autonomous learning.
-We conduct extensive experiments on the EMMA, Olympiad, and MathVista datasets, achieving promising results that outperform the current SOTA model by 15.84% across all tasks.
-Meanwhile, the additional analytical experiments also verify the model’s progress as well as generalization ability.
-1 Introduction
+Collaborative reasoning with multiple agents offers the potential for more robust and diverse problem-solving.
+However, existing approaches often suffer from homogeneous agent behaviors and lack of reflective and rethinking capabilities.
+We propose Multi-Agent Personality Shaping ((MAPS), a novel framework that enhances reasoning through agent diversity and internal critique.
+Inspired by the Big Five personality theory, MAPS assigns distinct personality traits to individual agents, shaping their reasoning styles and promoting heterogeneous collaboration.
+To enable deeper and more adaptive reasoning, MAPS introduces a Critic agent that reflects on intermediate outputs, revisits flawed steps, and guides iterative refinement.
+This integration of personality-driven agent design and structured collaboration improves both reasoning depth and flexibility.
+Empirical evaluations across three benchmarks demonstrate the strong performance of MAPS, with further analysis confirming its generalizability across different large language models and validating the benefits of multi-agent collaboration.
+Codehttps://github.com/exoskeletonzj/MAPS
+Introduction
+Solving complex reasoning problems (bhattacharya2024large; li2024multimodal; he2024olympiadbench) often requires more than just accurate perception and factual recall—it demands nuanced interpretation, multi-step inference, and the ability to adapt when initial attempts fail (fu2024isobench; li2024survey).
+While large language models (LLMs) demonstrate promising capabilities in solving such problems, they frequently fall short in scenarios requiring sustained reasoning, internal verification, and flexible strategy revision (anand2024mm; alasadi2024multimodal; gao2024cantor; wang2024t; xu2025large).
+A key challenge lies in how to effectively combine diverse reasoning strategies and incorporate mechanisms for intermediate reflection.
+Prior work has explored both single-agent solutions and simple collaborative setups (e.g., paired discussion or voting) (kaesberg2025voting), yet these approaches often suffer from rigid behaviors and limited capacity for self-correction.
+This raises an important research question: how can we enhance the depth and adaptability of reasoning by enabling more flexible and reflective solution processes?
+Figure 1 illustrates a representative scenario that embodies these challenges, where successful problem solving requires interpreting multimodal input and applying domain-specific reasoning.
 Figure 1: An example of a multimodal scientific multiple-choice problem. The correct answer is derived based on the reasoning over inputs that include context, question, and diagram.
-Multimodal scientific problems (MSPs) cover scientific scenarios that involve multiple modalities [5, 22, 16], such as text and vision.
-These problems typically span fields like mathematics, physics, and chemistry, requiring rigorous logical reasoning and solid domain expertise [3, 1, 13, 36, 41].
-In the realm of artificial intelligence, effectively addressing these cross-modal, multi-domain challenges remains both important and difficult [12, 20].
-Figure 1 illustrates a typical problem scenario that includes the context, the question statement, and an illustrative diagram.
-The diagram shows a lever divided into left and right sections: an unknown mass is hung on the left and a known mass on the right, each positioned at one-quarter of the total length from the respective ends, with a groove at the center.
-The accompanying description states that when the unknown mass is hung on the left, the system remains in static equilibrium.
-Solving this problem requires not only understanding the combined visual and textual information but also applying the lever balance principle by comparing the effects on both sides.
-One intuitive solution is to delegate the task-solving process to a single multimodal large language model (MLLM).
-Although the trending MLLMs have the basic abilities (e.g., diagram parsing and theorem retrieval), they are not well optimized to combine these skills in complex scenarios.
-Therefore, it motivates us to tackle this critical research question: how to leverage and elicit the off-the-shelf MLLMs to address the challenging MSPs?
-To address the problem, this work proposes to establish a multimodal multi-agent framework for multimodal scientific problem-solving.
-Motivated by the Big Seven Personality theory (i.e., Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity) [2, 4, 31], the MLLMs can function as independent agents with specific skills.
-Within the framework, these agents are guided to perform the problem-solving in a deliberate, cooperative manner.
-As one of the initial attempts for MSPs, we offer a novel solution for handling multimodal scientific scenarios.
-In solving MSPs, traditional approaches [35, 19, 15, 43, 6, 29] generally rely on a single MLLM to handle all tasks at once.
-Although these models possess basic abilities such as diagram parsing and theorem retrieval, they struggle to effectively integrate these skills in complex scenarios, resulting in what can be termed a multi-modal comprehensive reasoning challenge in scientific problem solving.
-This limitation arises because a single model cannot simultaneously manage diagram interpretation, context alignment, and professional knowledge supplementation, thus failing to mimic the step-by-step reasoning process of human cognition.
-In other words, while these models excel in specific tasks, they are unable to flexibly utilize various abilities to reason and solve complex multi-modal problems in a gradual, human-like manner.
-To address this, we propose a progressive four-agent solving strategy that sequentially employs a interpreter, an aligner, a scholar, and a solver, ultimately enhancing the accuracy and efficiency of solving multi-modal scientific problems.
-The lack of reflective and rethinking capabilities is another issue in traditional MSPs solving methods.
-Single MLLMs typically perform one-time reasoning and lack the ability for repeated reflection, adjustment, and iteration, which are essential in human problem-solving.
-When solving complex problems, humans continuously refine and adjust their solutions based on initial reasoning results, gradually approaching the correct answer.
-In contrast, traditional models often remain stuck at the initial reasoning stage, unable to self-correct and optimize.
-To address this, we designed a Critic agent, inspired by Socratic questioning [10, 27], which has unique advantages in fostering critical thinking and stimulating autonomous learning.
-Through a heuristic feedback mechanism, the Critic provides multiple rounds of feedback and continuous correction throughout the four-step solving process, simulating the iterative reflection process of humans.
-This design effectively overcomes the limitations of single models in terms of reflective and iterative capabilities, improving the flexibility and accuracy of the solution process.
-Figure 2: The corresponding relation between the Big Seven Personality theory and the seven function-specific agents.
-In summary, we propose a Multi-Agent framework for MSPs, based on the Big Seven Personality theory [2, 4, 31] and Socratic questioning (MAPS).
-This framework defines seven distinct roles—Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity—derived from the Big Seven theory, ensuring clear role delineation and mutual complementarity, as shown in Figure 2 and  Section 2.1.
-Together, these roles create a robust multi-agent architecture.
-Extensive experiments on the EMMA, Olympiad, and MathVista datasets show promising results across various reasoning benchmarks, confirming the effectiveness and superiority of our approach.
-Further analyses support our model’s advantages.
-The complete problem-solving process and data are provided in the supplementary materials.
-In summary, the contributions of this paper are threefold:
-$\bullet$ We propose a multi-agent framework based on the Big Seven personality theory.
-The framework integrates different personality traits from the Big Seven model to create seven complementary agents.
-To the best of our knowledge, this represents the first attempt in the field.
-$\bullet$ To tackle the challenges of multi-modal reasoning and iterative reflection, we propose a four-step strategy with a Critic agent inspired by Socratic questioning.
-This approach guides agents through structured stages, offering continuous feedback to refine reasoning.
-$\bullet$ Extensive experiments on various scientific domain datasets demonstrate the effectiveness and superiority of the MAPS framework, offering 15.84% performance gains.
-Additional analytical experiments further validate the interpretability of both the optimization process and its results.
-2 Methodology
-In this section, we provide a detailed introduction to MAPS.
-Our design is based on Large Language Models (LLMs) as intelligent agents capable of performing various functions.
-We define the task of solving MSPs as follows: Given an input diagram $d_{i}$, a context $c_{i}$, and a question $q_{i}$, the expected output is the answer $a_{i}$, which can be either an option or a computed result.
-The formal definition of this process is:
-$a_{i}=\mathcal{M}(d_{i},c_{i},q_{i}),$ (1)
-where $\mathcal{M}(d_{i},c_{i},q_{i})$ is the response obtained when the input diagram $d_{i}$, context $c_{i}$, and question $q_{i}$ are processed together by the model to generate the answer $a_{i}$.
-The goal of this task is to provide accurate answers through the model.
+Figure 2: Comparison of reasoning paradigms.
+Single-agent and two-agent approaches offer limited adaptability.
+MAPS enables dynamic collaborative reasoning.
+Right: Built on GPT-4o, MAPS achieves the best performance across three benchmarks.
+As illustrated in Figure 2, existing methods (wang2024examining; landau2024computational; hardiansyah2024analysis; zhang2024mm; caffagni2024revolution; qiu2024application) for complex reasoning problems often adopt single-agent solutions or simple two-agent collaborations.
+Although effective to some extent, these setups suffer from homogeneous agent behaviors—reasoning steps tend to repeat similar patterns, leading to redundancy and premature convergence.
+The lack of diversity limits exploration and reduces the chance of identifying alternative perspectives or correcting errors, especially in multi-turn settings where roles and strategies remain undifferentiated.
+Another issue is the lack of reflective and rethinking capabilities in existing approaches.
+As shown in Figure 2, the interaction between two agents is often linear and lacks mechanisms for feedback or revision.
+Even with multiple turns, agents rarely revisit earlier reasoning or correct initial misconceptions.
+In contrast, human reasoning is inherently iterative: People reflect, reassess, and adjust their thinking over time.
+Without structured reflection, current methods risk premature convergence and fail to recover from early-stage errors.
+To address these two issues, we propose MAPS (Multi-Agent Personality Shaping), a collaborative reasoning framework that enhances both diversity and adaptability in complex problem solving.
+Inspired by the Big Five personality theory (almagor1995big; benet1995big; simms2007big), MAPS shapes the reasoning behaviors of a set of role-specialized agents through distinct personality traits, promoting heterogeneous collaboration and mitigating behavioral homogeneity.
+To enable reflective thinking and iterative refinement, MAPS further introduces a Critic agent that revisits intermediate outputs, identifies flawed steps, and provides structured feedback.
+This integration of personality-guided agent design and internal critique supports deeper, more flexible reasoning aligned with human cognitive processes.
+As shown in Figure 3, the interpreter (Openness) explores information from multiple angles, the aligner (Agreeableness) reconciles visual and textual cues, the scholar (Conscientiousness) ensures precision and factual grounding, the solver (Extraversion) drives goal-oriented conclusions, and the critic (Neuroticism) questions assumptions and detects flaws.
+The critic is further inspired by Socratic questioning (elder1998role; paul2019thinker), providing reflective feedback throughout the multi-stage reasoning process.
+Together, these roles enable a structured yet flexible collaboration that supports deeper and more reliable problem solving.
+We conduct extensive experiments on three challenging benchmarks: MathVista (lu2023mathvista), OlympiadBench (he2024olympiadbench), and EMMA (hao2025can).
+These datasets cover a wide range of complex reasoning tasks.
+MAPS consistently outperforms baseline methods across all datasets, confirming its effectiveness in enhancing both reasoning depth and adaptability.
+We further evaluate MAPS under different base LLMs, and results consistently show its superiority across model backbones.
+Additional analyses examine the impact of the feedback mechanism and the overall efficiency of the framework.
+Our main contributions are as follows:
+$\bullet$ We propose MAPS, a multi-agent reasoning framework.
+To the best of our knowledge, this is the first work that incorporates personality shaping based on the Big Five theory into collaborative reasoning.
+$\bullet$ MAPS addresses two key challenges in existing methods: behavioral homogeneity and lack of reflection.
+It assigns distinct personality traits to agents and introduces a Critic inspired by Socratic questioning.
+$\bullet$ We conduct extensive experiments on three scientific reasoning benchmarks.
+MAPS achieves consistent performance gains (up to 15.84%) and generalizes well across different tasks and base models.
+Figure 3: The corresponding relation between the Big Five Personality theory and the five function-specific agents.
+Methodology
 Figure 4: The overall architecture of MAPS.
-It illustrates seven functional agents based on the Big Seven Personality theory.
-It first includes the Manager agent with predefined interaction logic and the UserProxy agent responsible for receiving user inputs.
-Subsequently, four specialized agents—Interpreter, Aligner, Scholar, and Solver—are introduced, each corresponding to a specific step in solving MSPs.
-Finally, the Critic agent is presented, providing feedback and corrections to ensure the results are more accurate and interpretable.
-The model architecture for solving MSPs is shown in Figure 3.
-In Section 2.1, we introduce the seven distinct functional agents designed based on the Big Seven Personality theory.
-In Section 2.2, we present the predefined interaction logic in the Manager agent, along with the UserProxy agent responsible for receiving user input.
-In Section 2.3, we discuss the four specific agents—Interpreter, Aligner, Scholar, and Solver——each representing a step in solving MSPs.
-Finally, in Section 2.4, we introduce the Critic agent, which provides feedback and corrections for the specific solution steps to ensure the results are more accurate and interpretable.
-2.1 Big Seven Personality
-The Big Seven Personality theory is an extension of the traditional Big Five Personality [33, 8, 7], adding two other dimensions: Self-Esteem and Sensitivity.
-It divides human personality into seven main dimensions: Conscientiousness, Agreeableness, Extraversion, Neuroticism, Openness, Self-Esteem, and Sensitivity.
-Each reflects unique tendencies in emotions, cognition, and behavior, revealing the diversity among individuals in decision-making, communication, and problem-solving.
-This theory provides a systematic framework for understanding individual differences and offers valuable theoretical support and guidance for psychological research and practical applications.
-Inspired by the Big Seven Personality theory, we designed a multi-agent system that strives to maintain both diversity and integrity in completing complex tasks.
-As shown in Figure 2, the seven personality traits correspond to seven agents, each performing different roles.
-The Manager agent represents Conscientiousness, responsible for creating the experimental plan and schedule, ensuring each step is executed according to the predefined plan.
-The UserProxy agent represents Agreeableness, ensuring smooth information flow and coordinating the allocation of tasks within the experiment.
-The Interpreter agent represents Extraversion, responsible for interpreting diagrams into captions, thereby providing new ideas and information for MSPs to foster innovation.
-The Aligner agent represents Neuroticism, tasked with aligning the caption, context, and question to ensure the safe integration of these elements.
-The Scholar agent represents Openness, responsible for researching the professional knowledge required by MSPs and exploring various hypotheses.
-The Solver agent represents Self-Esteem, in charge of gathering all necessary information and resolving MSPs by selecting the most appropriate experimental approach.
-The Critic agent represents Sensitivity, providing feedback on the four solution steps performed by the Interpreter, Aligner, Scholar, and Solver agents, and offering evaluations along with adjustment strategies.
-The specific duties and role distribution of these agents are discussed in detail in the following sections.
-2.2 Agents Interaction Logic
-This section primarily introduces the interaction logic of the seven personality agents.
-Since solving MSPs requires multimodal semantic integration and multi-step reasoning within specialized domains, the Manager agent is responsible for formulating a detailed research plan and interaction logic to ensure the smooth execution of various system tasks.
-As shown in Figure 3, the Manager agent coordinates the operations of all other agents to ensure the task is completed successfully.
-Additionally, the UserProxy agent handles receiving user input, ensuring smooth communication between the user and the system, and facilitating effective interaction.
-Algorithm 1 MAPS Procedure
-1:  Input: Diagram $d_{i}$, Context $c_{i}$, and Question $q_{i}$
-2:  Output: Final answer $a_{i}$
-3:  Manager: Formulate a detailed research plan
-4:  UserProxy: Receive the input $x\in$ Dataset $D$
-5:  Interpreter: Interpret the diagram into a caption $p_{i}=\mathcal{M}_{\text{int}}(d_{i})$
-6:  Aligner: Align the caption with the question and context $l_{i}=\mathcal{M}_{\text{ali}}(p_{i},c_{i},q_{i})$
-7:  Scholar: Research for MSPs $s_{i}=\mathcal{M}_{\text{sch}}(l_{i},p_{i},c_{i},q_{i})$
-8:  Solver: Solve based on the complete information $a_{i}=\mathcal{M}_{\text{sol}}(s_{i},l_{i},p_{i})$
-9:  Critic: Score and assess the four steps $r_{i}=\mathcal{M}_{\text{crit}}(a_{i},s_{i},l_{i},p_{i})$
-10:  while True do
-11:     if All Scores are 5 then Break
-12:     else Choose the minimum score from $r_{i}$ and return to the step $\in\{\textit{Interpreter},\textit{Aligner},\textit{Scholar},\textit{Solver}\}$ to optimize the process.
-13:     end if
-14:  end while
-15:  return  Final answer $a_{i}$
-As illustrated in Algorithm 1, the entire workflow is orchestrated by the Manager agent, corresponding to the flowchart shown in Figure 3.
-Once the UserProxy agent receives the user input, the system proceeds through a four-step problem-solving process involving the Interpreter, Aligner, Scholar, and Solver agents.
-After completing these steps, the Critic agent, based on Socratic questioning, evaluates and scores the results of each step.
-2.3 Four-Step MSPs Solving
-This section will provide a detailed introduction to the functions of the Interpreter, Aligner, Scholar, and Solver agents.
+The framework consists of five functional agents inspired by the Big Five personality theory.
+The core reasoning process is carried out by four specialized agents—Interpreter, Aligner, Scholar, and Solver—each responsible for a distinct stage in solving complex reasoning problems.
+Finally, the Critic agent provides reflective feedback and correction to enhance accuracy and interpretability.
+This section introduces MAPS in four key components: Preliminaries, Agentic Interaction Logic, the Four-Step Reasoning process, and the Critic and Feedback mechanism.
+Preliminaries
+Task Definition.
+We define the complex reasoning task as a mapping from structured multimodal inputs to a target answer space.
+Let $\mathcal{D}$, $\mathcal{C}$, and $\mathcal{Q}$ denote the input spaces of diagrams, contexts, and questions, respectively, and let $\mathcal{A}$ denote the answer space.
+Each instance is a triplet $(d_{i},c_{i},q_{i})\in\mathcal{D}\times\mathcal{C}\times\mathcal{Q}$, and the goal is to predict the corresponding answer $a_{i}\in\mathcal{A}$.
+The task is thus modeled as a function:
+$\mathcal{M}:\mathcal{D}\times\mathcal{C}\times\mathcal{Q}\rightarrow\mathcal{A},\quad a_{i}=\mathcal{M}(d_{i},c_{i},q_{i}),$ (1)
+where $\mathcal{M}$ denotes the reasoning system responsible for processing visual-textual input and generating the output.
+Agent-Based Modeling.
+In MAPS, the reasoning process $\mathcal{M}$ is decomposed into a sequence of collaborative stages executed by a set of role-specialized agents $\{\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{K}\}$, where each agent $\mathcal{A}_{k}$ performs a specific function conditioned on both the original input and intermediate reasoning states.
+Let $\mathcal{S}_{0}=(d_{i},c_{i},q_{i})$ be the initial input state. The overall reasoning unfolds as a staged transformation:
+$\mathcal{S}_{k}=\mathcal{A}_{k}(\mathcal{S}_{k-1}),\quad\text{for }k=1,2,\dots,K,$ (2)
+where $\mathcal{S}_{k}$ denotes the intermediate reasoning state after the $k$-th agent’s operation.
+The final answer is extracted as $a_{i}=\mathrm{Extract}(\mathcal{S}_{K})$.
+This formulation reflects the staged reasoning in MAPS, where specialized agents collaboratively refine the solution.
+Agents Interaction Logic
+As illustrated in Algorithm 1, this section introduces the interaction logic among the five personality-driven agents in MAPS.
+Complex problem reasoning requires multimodal semantic integration and structured inference over multiple steps.
+MAPS models this process as a functional composition of role-specialized reasoning agents, each shaped by a distinct personality trait.
+Let $\mathbf{x}=(d_{i},c_{i},q_{i})$ be the input, and let $\mathbf{p}_{k}\in\mathbb{R}^{m}$ denote the personality embedding associated with agent $\mathcal{A}_{k}$.
+The entire collaborative reasoning process is represented as:
+$a_{i}=\mathcal{F}(\mathbf{x};\mathbf{p}_{1},\dots,\mathbf{p}_{4})=\mathcal{A}_{4}\circ\mathcal{A}_{3}\circ\mathcal{A}_{2}\circ\mathcal{A}_{1}(\mathbf{x}),$ (3)
+where each agent $\mathcal{A}_{k}(\cdot;\mathbf{p}_{k})$ executes its stage conditioned on its personality vector $\mathbf{p}_{k}$, contributing a unique reasoning perspective.
+After the initial reasoning trajectory $\mathcal{T}=\{\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4}\}$ is generated, the Critic agent performs a reflective evaluation by applying a feedback function:
+$\mathbf{f}=\texttt{Reflect}(\mathcal{T}),\quad\mathbf{f}=\{f_{1},f_{2},f_{3},f_{4}\}\in\mathbb{R}^{4},$ (4)
+where each score $f_{k}$ reflects the confidence of the $k$-th stage.
+If any $f_{k}<\tau$, the corresponding stage is revised.
+This feedback loop complements the forward reasoning path, allowing MAPS to emulate deliberative human cognition through structured collaboration and critique.
+Proposition 1 (Monotonic Free-Energy Descent).
+For the free energy $F^{(t)}=\mathbb{E}_{q^{(t)}}[-\!\log p(\mathbf{x},a_{i}\mid\theta)]+\mathrm{KL}\bigl(q^{(t)}\|\;p\bigr)$, each Critic-triggered update satisfies
+$F^{(t+1)}\;\leq\;F^{(t)}.$ (5)
+Thus the MAPS iteration produces a non-increasing free-energy sequence that converges to a stationary point. The full proof is provided in Appendix A.1.
+Algorithm 1 MAPS Collaborative Reasoning with Reflective Feedback
+1:Input: Diagram $d_{i}$, Context $c_{i}$, Question $q_{i}$, Personality vectors $\{\mathbf{p}_{1},\dots,\mathbf{p}_{4}\}$
+2:Initialize: $\mathcal{S}_{0}=(d_{i},c_{i},q_{i}),\quad t\leftarrow 0$
+3:repeat
+4:  Interpreter:  $p_{i}=\mathcal{A}_{1}(\mathcal{S}_{0};\mathbf{p}_{1})$
+5:  Aligner:  $l_{i}=\mathcal{A}_{2}(p_{i},c_{i},q_{i};\mathbf{p}_{2})$
+6:  Scholar:  $s_{i}=\mathcal{A}_{3}(l_{i},p_{i},c_{i},q_{i};\mathbf{p}_{3})$
+7:  Solver:  $a_{i}=\mathcal{A}_{4}(s_{i},l_{i},p_{i};\mathbf{p}_{4})$
+[…]
+9:  Critic:  $\mathbf{f}^{(t)}=\texttt{Reflect}(\mathcal{T}^{(t)}),\quad f_{k}\in[0,1]$
+10:  if all $f_{k}\geq\tau$ then
+11:   break
+12:  else
+13:   Identify stage $k^{*}=\arg\min_{k}f_{k}$
+14:   Rerun agent $\mathcal{A}_{k^{*}}$ with updated input
+15:   $t\leftarrow t+1$
+16:  end if
+17:until convergence
+18:return Final answer $a_{i}$
+Four-Step Reasoning
+Given input $\mathbf{x}=(d_{i},c_{i},q_{i})$, MAPS conducts a structured reasoning process across four personality-driven agents: Interpreter, Aligner, Scholar, and Solver.
+Each agent $\mathcal{A}_{k}$ is parameterized by a personality embedding $\mathbf{p}_{k}\in\mathbb{R}^{m}$, shaping its reasoning style and focus.
+The entire inference pipeline can be expressed as a function composition:
+$a_{i}=\mathcal{A}_{4}\circ\mathcal{A}_{3}\circ\mathcal{A}_{2}\circ\mathcal{A}_{1}(\mathbf{x};\mathbf{p}_{1},\mathbf{p}_{2},\mathbf{p}_{3},\mathbf{p}_{4}),$ (6)
+where each intermediate state $\mathcal{S}_{k}$ is defined recursively as $\mathcal{S}_{k}=\mathcal{A}_{k}(\mathcal{S}_{k-1};\mathbf{p}_{k})$, with $\mathcal{S}_{0}=\mathbf{x}$.
 Interpreter.
-To effectively integrate visual context into the textual context, we adopt a multi-modal fusion paradigm using the Interpreter agent.
-Initially, the Interpreter agent is employed to analyze diagrams meticulously, converting them into detailed and accurate captions.
-This process ensures that rich visual details are faithfully captured in the textual description, such as relative positions, dimensions, and other features.
-Subsequently, the generated caption is fused with the textual context and the question in later steps, achieving a cohesive integration of information that provides a richer and more consistent basis for problem-solving.
-The formula is shown as follows:
-$p_{i}=\mathcal{M}_{\text{int}}(d_{i}),$ (2)
-where $\mathcal{M}_{\text{int}}(\cdot)$ represents the Interpreter agent and $p_{i}$ denotes the caption.
+The Interpreter agent aims to extract structured visual semantics from diagram $d_{i}$, translating them into a caption representation $p_{i}$ that can be consumed by downstream language agents.
+Let $\phi_{\text{vis}}(d_{i})$ be a visual encoder, and $\psi_{\text{lang}}(\cdot)$ a caption generator.
+Then the agent performs:
+$p_{i}=\mathcal{A}_{1}(d_{i};\mathbf{p}_{1})=\psi_{\text{lang}}\big(\phi_{\text{vis}}(d_{i})+W_{1}\mathbf{p}_{1}\big),$ (7)
+where $W_{1}\in\mathbb{R}^{d\times m}$ projects the personality embedding into the encoder space, modulating its attention to attributes like spatiality, color, or topology.
 Aligner.
-In textual form, the information conveyed by elements such as the context, questions, and options differs from the content presented in the diagram.
-These differences can lead to discrepancies in understanding the problem, which in turn may affect subsequent reasoning and judgment.
-Therefore, in order to gain a deeper understanding of the problem statement and accurately capture multimodal semantic information, precise alignment of these information granularities is crucial.
-As shown below, $l_{i}$ illustrates the result of the alignment between the diagram, context, and question:
-$l_{i}=\mathcal{M}_{\text{ali}}(p_{i},c_{i},q_{i}),$ (3)
-where $\mathcal{M}_{\text{ali}}(\cdot)$ represents the aligner agent.
+To resolve semantic mismatches across modalities, the Aligner fuses the interpreted diagram caption $p_{i}$ with the textual context $c_{i}$ and question $q_{i}$.
+The process outputs an alignment representation $l_{i}$, optimized to preserve shared semantics and suppress modality conflict.
+Formally,
+$l_{i}=\mathcal{A}_{2}(p_{i},c_{i},q_{i};\mathbf{p}_{2})=\mathrm{CrossFuse}\big(p_{i},c_{i},q_{i};\mathbf{p}_{2}\big),$ (8)
+where $\mathrm{CrossFuse}(\cdot)$ denotes a multi-head attention-based fusion operator, adaptively weighted by $\mathbf{p}_{2}$ to emphasize visual or linguistic cues based on agent bias.
+Scholar.
+While $l_{i}$ captures semantic consistency, complex reasoning often requires external knowledge supplementation.
+The Scholar agent retrieves and integrates domain-specific knowledge $\mathcal{K}(l_{i})$, such as physics principles or mathematical theorems.
+We define:
+$s_{i}=\mathcal{A}_{3}(l_{i},p_{i},c_{i},q_{i};\mathbf{p}_{3})=\mathrm{KnowAug}\big(l_{i},\mathcal{K}(l_{i});\mathbf{p}_{3}\big),$ (9)
+where $\mathrm{KnowAug}(\cdot)$ augments contextual embeddings with retrieved tuples $\mathcal{K}(l_{i})$ from a structured knowledge memory, and $\mathbf{p}_{3}$ biases the agent toward formal rigor or heuristic reasoning.
+Solver.
+The Solver agent aggregates all upstream outputs and executes logical composition to generate the final answer $a_{i}$.
+Let $\mathcal{H}_{i}=\{p_{i},l_{i},s_{i}\}$ be the hybrid reasoning state.
+The solver computes:
+$a_{i}=\mathcal{A}_{4}(\mathcal{H}_{i};\mathbf{p}_{4})=\mathrm{Deduct}\big(p_{i},l_{i},s_{i};\mathbf{p}_{4}\big),$ (10)
+where $\mathrm{Deduct}(\cdot)$ is a constrained generation module that synthesizes the inputs under logical, numerical, or symbolic rules.
+The final prediction $a_{i}\in\mathcal{A}$ may be a selected option or a free-form answer.
+Together, these four stages construct a trajectory $\mathcal{T}=\{p_{i},l_{i},s_{i},a_{i}\}$, on which the Critic agent operates for reflection and feedback.
+Proposition 2 (Collaborative Information Bottleneck).
+Let $\mathcal{S}_{k}$ be the intermediate output of the $k$-th agent, given input $\mathbf{x}=(d_{i},c_{i},q_{i})$ and target answer $a_{i}$.
+Then the MAPS reasoning process optimizes
+$\min\sum_{k=1}^{4}I(\mathbf{x};\mathcal{S}_{k})\quad\text{s.t.}\quad I(\mathcal{S}_{k};a_{i})\geq\varepsilon,$ (11)
+where $I(\cdot;\cdot)$ denotes mutual information. The Critic monitors constraint violations and reactivates stages with insufficient task-relevant information.
+A derivation and discussion are provided in Appendix A.2.
 Models CoT MathVista EMMA OlympiadBench Avg.
 Gen. Math Avg. Math Phy. Chem. Avg. MECO MZCE MZCO PECO PZCE Avg.
-Random Choice - 26.09 22.78 24.30 13.00 23.00 27.00 23.00 0.67 0.33 0.00 1.75 0.33 0.87 16.06
-Human Expert - 56.09 55.74 55.90 75.00 64.50 86.00 64.50 48.00 34.67 30.36 54.17 12.33 37.80 52.73
+Random Choice - 26.09 22.78 24.30 13.00 23.00 27.00 21.00 0.67 0.33 0.00 1.75 0.33 0.87 16.06
+Human Expert - 56.09 55.74 55.90 75.00 64.50 86.00 75.17 48.00 34.67 30.36 54.17 12.33 37.80 52.73
 Claude 3.5 Sonnet - 68.04 63.15 65.40 23.00 34.00 44.00 33.67 20.67 13.00 10.71 10.75 14.00 13.23 37.43
@@
 LLaVA-Onevision-72B - 62.83 58.52 60.50 25.00 32.00 24.00 27.00 6.67 7.33 3.57 3.29 9.67 6.18 31.23
-[…]
+Claude 3.5 Sonnet 71.74 64.26 67.70 30.00 38.00 41.00 36.33 24.00 11.00 16.07 12.72 10.33 13.23 39.09
 Gemini 2.0 Flash 70.22 75.56 73.10 24.00 41.00 36.00 33.67 12.67 6.33 3.57 4.61 2.33 5.39 37.38
@@
 MAPS (GPT-4obase) - 75.87 83.15 79.80 52.00 71.00 51.00 58.00 46.00 30.33 32.14 28.51 28.33 31.14 56.31
-Table 1: Performance of different models across 10 subtasks from 3 datasets.
-‘Gen.’ refers to General problems in the Mathvista dataset, ‘Phy.’ and ‘Chem.’ denote Physics and Chemistry tasks in the EMMA dataset.
-‘MECO’, ‘MZCO’, and ‘MZCE’ correspond to COMP problems in English, COMP problems in Chinese, and CEE problems in Chinese, respectively, within the OlympiadBench dataset’s mathematics tasks.
-‘PECO’ and ‘PZCE’ represent the COMP problems in English and CEE problems in Chinese under the physics tasks of the OlympiadBench dataset.
-Scholar.
-After obtaining the semantically fused $l_{i}$, the model has essentially grasped the main idea of the problem.
-However, for scientific questions, mere comprehension is insufficient, as they typically require profound domain-specific knowledge, which is crucial for problem-solving.
-As illustrated in Figure 1, certain problems necessitate mastery of specialized formulas, such as those related to the principles of lever balance.
-Therefore, the Scholar agent is responsible for acquiring and integrating domain-specific knowledge based on $l_{i}$, thereby providing the necessary theoretical and practical support for the final solution $s_{i}$:
-$s_{i}=\mathcal{M}_{\text{sch}}(l_{i},p_{i},c_{i},q_{i}),$ (4)
-where $\mathcal{M}_{\text{sch}}(\cdot)$ represents the Scholar agent.
-Solver.
-After the three steps involving the Interpreter, Aligner, and Scholar agents, the system has extracted and fused extensive information from both visual and textual modalities, providing rich contextual and specialized knowledge for the task.
-However, the real challenge lies in synthesizing this fragmented information into a coherent and logically sound solution.
-The fourth step, carried out by the Solver agent, bears the greatest pressure.
-This agent must not only integrate the information gathered in previous steps but also engage in deep reasoning to derive the final answer.
-With access to the most comprehensive data, the Solver has a unique advantage, allowing it to simplify complex issues and ensure the final answer $a_{i}$ is accurate and meets the expected standards, as given by:
-$a_{i}=\mathcal{M}_{\text{sol}}(s_{i},l_{i},p_{i}),$ (5)
-where $\mathcal{M}_{\text{sol}}(\cdot)$ represents the Solver agent.
-2.4 Critic and Feedback
-Critic.
-This section introduces the Critic agent, which is responsible for evaluating and providing feedback on the four-step MSPs solving process.
-Rooted in Socratic questioning, the Critic prompts agents to reflect on their reasoning by asking questions such as “What assumptions are you making?” or “How can you justify this decision?” This approach encourages critical thinking and self-examination, key aspects of the Socratic method.
-Rather than relying on the final solution labels, the Critic evaluates the logic and justification of each step.
-The goal is to uncover flaws and push agents to refine their thought processes.
-Based on this feedback, the system identifies the weakest step and initiates a rollback and a redo to enhance the problem-solving process.
-The formula is given by:
-$r_{i}=\mathcal{M}_{\text{crit}}(a_{i},s_{i},l_{i},p_{i}),$ (6)
-where $\mathcal{M}_{\text{crit}}(\cdot)$ represents the Critic agent. $r_{i}$ is the scores and assessments of the four steps.
-3 Experiments
-We conduct extensive experiments on three MSPs datasets.
-This section introduces the three MSPs datasets and baselines in Section 3.1, and presents the experimental results and analysis in Section 3.2.
-The analysis and statistics of the Critic agent are presented in Section 3.3.
-3.1 Datasets and Baselines
-This study uses the latest three MSPs datasets: MathVista [23], OlympiadBench [16], and EMMA [14].
-MathVista focuses on mathematics and general domains to assess models’ ability to solve complex scientific problems; OlympiadBench targets mathematics and physics to evaluate performance on Olympiad-level challenges; and EMMA, a multimodal dataset covering mathematics, physics, and chemistry, tests interdisciplinary scientific reasoning.
-Detailed descriptions of the tasks and datasets are in Appendix A, along with domain abbreviations, full names, and dataset descriptions in Table 1.
-In the experimental setup, we select GPT-4o [^2] https://openai.com/api/ as our agent.
-For the baselines, we incorporate the current mainstream MLLMs, both closed-source and open-source, and evaluated them in both direct and CoT settings.
-The experimental settings and baselines are detailed in Appendix B.
-3.2 Main Results
-MAPS has set a new state-of-the-art (SOTA) performance, surpassing human-level performance for the first time.
-The results in Table 1 provide a detailed comparison of MAPS and the baselines.
-As shown, MAPS outperforms the current SOTA by 15.84% across all tasks, demonstrating its superiority in MSPs problem-solving.
-Additionally, MAPS slightly exceeds human experts by 3.58% in overall performance.
-MAPS excels in mathematical, physical, chemical, and general problems, demonstrating strong interdisciplinary reasoning.
-The multi-agent system, representing the Big Seven personalities, sets a new SOTA in collaboration.
-Experimental results show MAPS’s advantage in multimodal semantic fusion and multi-step reasoning, leveraging its ability to integrate diagrams with context and questions.
-This boosts problem-solving by enhancing domain-specific knowledge.
-The Critic agent, using Socratic feedback, offers suggestions for improvement, improving accuracy and reliability in complex tasks.
+Table 1: Performance across 10 subtasks.
+Gen.
+= General (MathVista), Phy./Chem.
+= Physics/Chemistry (EMMA), MECO/ MZCO/ MZCE = English/Chinese COMP & CEE Math (OlympiadBench), PECO/ PZCE = English/Chinese Physics (OlympiadBench).
+Critic and Feedback
+The Critic agent evaluates the internal reasoning trajectory $\mathcal{T}=\{p_{i},l_{i},s_{i},a_{i}\}$ without relying on ground-truth answers.
+Inspired by Socratic questioning, it examines each stage’s logic and justification to identify flawed assumptions and incomplete inferences.
+We define the feedback vector as:
+$\mathbf{f}_{i}=\mathcal{M}_{\text{crit}}(p_{i},l_{i},s_{i},a_{i}),\quad\mathbf{f}_{i}\in[0,1]^{4},$ (12)
+where each element $f_{i}^{(k)}$ represents the Critic’s confidence in the correctness and completeness of stage $k$.
+The weakest stage is selected by:
+$k^{\ast}=\arg\min_{k}f_{i}^{(k)},\quad\text{if}\quad f_{i}^{(k^{\ast})}<\tau,$ (13)
+which triggers a targeted revision. This reflection-driven loop promotes iterative self-correction and deepens reasoning reliability.
+Experiments
+Datasets and Baselines
+We evaluate on three benchmarks for complex problem reasoning: MathVista, OlympiadBench, and EMMA.
+MathVista covers math and general science, OlympiadBench focuses on high-level math and physics, and EMMA includes math, physics, and chemistry.
+See Appendix B for details.
+We use GPT-4o (achiam2023gpt) as the base model.
+For comparison, we include leading multimodal large language models (MLLMs), both proprietary and open-source, tested under both direct and Chain-of-Thought (CoT) settings.
+Experimental details and baseline setups are provided in Appendix C.
+Main Results
+MAPS achieves a new state-of-the-art (SOTA), surpassing human-level performance for the first time.
+As shown in Table 1, MAPS outperforms previous SOTA models by 15.84% and exceeds human expert performance by 3.58% across all tasks, highlighting its strength in solving complex multimodal problems.
+MAPS demonstrates robust performance across mathematical, physical, chemical, and general tasks, showcasing strong interdisciplinary reasoning.
+Its multi-agent design, based on the Big Five personality theory, enables effective collaboration and contributes to the SOTA results.
+The system excels in multimodal semantic integration and multi-step reasoning by jointly leveraging diagrams, contexts, and questions.
+Furthermore, the Critic agent applies Socratic feedback to refine responses, enhancing both accuracy and reliability on challenging tasks.
+MAPS exhibits strong adaptability and robustness across diverse reasoning tasks.
+The evaluation datasets span a wide range of question types, modalities, and reasoning difficulties.
+MathVista features judgment, multiple-choice, and open-ended fill-in-the-blank questions with varied answer formats, requiring accurate intent understanding and response generation.
+OlympiadBench emphasizes challenging open-ended problems demanding multi-step symbolic reasoning, where small errors can lead to divergent outcomes.
+EMMA introduces multimodal complexity with diagrams embedded in both questions and answer choices.
+Through feedback-driven multi-agent collaboration and Socratic questioning, MAPS effectively handles these challenges, achieving SOTA and demonstrating strong generalization across heterogeneous reasoning scenarios.
+Analysis of Critic Agent
+Critic schema includes scoring and Socratic feedback.
+Each reasoning step is rated from 0–5, guiding whether the system should backtrack or proceed.
+Feedback is heuristic, encouraging rethinking rather than offering direct answers.
+The Critic uses these scores to decide whether to regenerate specific steps, ensuring robustness in the final answer.
+The Critic enhances reasoning via Socratic feedback without using gold labels.
+As shown in Figure 5, the top illustrates the feedback schema, while the bottom shows feedback proportions across three datasets.
+The Critic prompts reflection and correction by encouraging agents to question assumptions rather than passively accept reasoning.
+Solver receives the most feedback in EMMA and OlympiadBench.
+As shown in the lower half of Figure 5, feedback varies by dataset.
+For MathVista, most steps need no regeneration, aligning with our superior 5.0% SOTA improvement in Table 1.
+This reflects strong baseline reasoning.
+In contrast, EMMA and OlympiadBench show the highest feedback for the Solver, especially in the interpretation, alignment, and integration steps.
+These are the most complex and error-prone stages.
+Other agents receive comparable and lower feedback, indicating better relative performance in their sub-tasks.
 Figure 5: The schema of the Critic agent, as well as the feedback and backtracking situations of the Critic agent across different datasets.
-In the MathVista dataset, the problem types are diverse, covering judgment questions, multiple-choice questions, and open-ended fill-in-the-blank questions, with various answer types such as text, floats, and integers.
-This requires the system to possess specialized capabilities, enabling it to accurately understand the question’s intent and provide specific responses.
-In well-known open-ended problems like those in OlympiadBench, the solution process involves complex steps and high difficulty, with open-ended answers that can lead to completely different outcomes due to subtle changes.
-In the EMMA dataset, the questions not only involve diagrams, but some options also contain diagrams, significantly increasing the challenge of understanding and aligning multiple diagrams.
-Furthermore, the question types are equally diverse, including open-ended and multiple-choice questions, raising the demands on the solving methods.
-The MAPS framework, through feedback-based multi-agent collaborative learning and the use of Socratic questioning, successfully addresses these different types of sub-tasks, achieving the latest SOTA.
-It demonstrates MAPS’s perfect adaptability and robustness when dealing with the diversity of question types, task difficulties, and multi-step reasoning.
-3.3 Analysis of Critic Agent
-The Critic agent plays a crucial feedback role in MAPS, enhancing the overall reasoning process without the reliance on gold labels.
-As shown in Figure 4, the upper part introduces the critic and feedback schema, while the lower part presents the feedback proportions provided by the Critic agent to each of the agents across three datasets, visually illustrating the feedback and corrections made by each assistant agent.
-The Critic agent’s feedback mechanism is rooted in Socratic questioning, which encourages deep reflection and rethinking of the reasoning steps.
-Instead of simply evaluating the process, the Critic fosters critical thinking, prompting the system to question its assumptions and improve the approach.
-The schema of the Critic agent consists of two components: scores and feedback.
-The scores record the ratings given by the Critic agent for the four problem-solving steps, ranging from 0 to 5.
-Higher scores indicate better execution, while lower scores highlight areas needing improvement.
-Feedback primarily captures the Critic agent’s Socratic-style, heuristic suggestions designed to encourage deeper reflection on each step.
-These suggestions are not limited to evaluations; they also guide the agents to reconsider and correct their reasoning process.
-For example, the Critic may prompt the agents to think about alternative approaches or reassess their conclusions, challenging them to think more critically.
-To ensure the final result’s quality, a rule is set in the Manager agent: if any step receives a low score, the system will backtrack and make corrections.
-If all steps score well, the system outputs the result directly.
-This approach guarantees the effectiveness of each step, optimizing the overall problem-solving process.
-In the EMMA and OlympiadBench datasets, the Solver agent exhibits the highest feedback proportion.
-The lower part of Figure 4 shows the distribution of agents that performed feedback corrections across the three datasets.
-It can be observed that the feedback proportions vary across the three datasets.
-In the MathVista dataset, the highest feedback proportion corresponds to no regeneration, indicating that the overall solution performance on this dataset is still quite good.
-This result is consistent with the findings in Table 1, and we have surpassed the previous SOTA by a 5.0% improvement.
-This suggests that while there is still room for further enhancement in certain areas, the system’s overall reasoning capability is already quite impressive when dealing with more complex problems.
-In contrast, in the EMMA and OlympiadBench datasets, the Solver agent shows the highest feedback proportion, suggesting that this part is the most difficult and prone to errors in these datasets.
-Specifically, this part includes the first three steps: interpretation, alignment, and integration of information provided by the Scholar.
-These steps require the integration of a large amount of data, involve the most complex reasoning, and bear the most critical responsibility—providing a complete and accurate solution to the problem.
-In comparison, the feedback proportions from the other three agents are similar, indicating that their error rates are relatively low, and the performance of these sub-steps is relatively better.
-4 Supplementary Analysis
-To conduct more comprehensive experiments, this section covers the following topics: ablation experiments in Section 4.1, the base model generalization in Section 4.2, the time efficiency of MSPs solving in Section 4.3.
-Due to space limitations, generalization experiments on other datasets, such as DiagramQG, are discussed in Appendix C.
-The case study and complete process for solving the MSPs task is outlined in Appendix D, and the prompts for all agents are provided in Appendix E.
-4.1 Ablation Study
+Supplementary Analysis
+Due to space limitations, generalization experiments on other datasets, such as DiagramQG (zhang2025diagram), are discussed in Appendix D.
+The case study and complete process of MAPS is outlined in Appendix E, and the prompts for agents are provided in Appendix F.
+Ablation Studies
+Ablating the Interpreter results in the greatest loss of performance.
+We conduct ablation experiments on the OlympiadBench dataset to evaluate the impact of each module on the overall performance.
+Table 2 presents the effects of removing the Interpreter, Aligner, Scholar, and Critic modules from the MAPS framework.
+The results show that removing the Interpreter agent causes the largest performance degradation, at 16.09%.
+This is because, in complex problem reasoning, diagrams contain a wealth of valuable information, which serves as an important supplement to the text.
+Understanding diagrams plays a crucial role in problem-solving.
+The removal of the Critic agent causes the smallest performance loss.
+It results in only a 7.05% decrease, underscoring its role in providing feedback and corrections.
+While this mechanism allows MAPS to backtrack and refine its reasoning, its impact is less than that of other agents.
+Removing the Scholar agent results in 11.49% performance drops, highlighting the importance of searching and integrating domain-specific knowledge.
+Finally, the removal of the Aligner agent causes a 10.86% drop, indicating that while diagram and context alignment is valuable, its effect is smaller compared to other components.
 Variation MECO MZCE MZCO PECO PZCE Avg.
@@
 We perform ablation experiments on the solving module ${w/o}_{\text{Interpreter}}$, ${w/o}_{\text{Aligner}}$, ${w/o}_{\text{Scholar}}$ or ${w/o}_{\text{Critic}}$ modules to evaluate the impact of removing these components.
-Ablating the Interpreter results in the greatest loss of performance.
-We conduct ablation experiments on the OlympiadBench dataset to evaluate the impact of each module on the overall performance.
-Table 2 presents the effects of removing the Interpreter, Aligner, Scholar, and Critic modules from the MAPS framework.
-The results show that removing the Interpreter agent causes the largest performance degradation, at 16.09%.
-This is because, in MSPs tasks, diagrams contain a wealth of valuable information, which serves as an important supplement to the text.
-Understanding diagrams plays a crucial role in problem-solving.
+Base Model Generalization
+MAPS improves performance across diverse base models.
+We conduct experiments to verify whether our MAPS framework demonstrates robust generalization across various base LLMs.
+The results confirm MAPS’s robustness and transferability, highlighting its adaptability and consistent performance across different foundation models.
+To further validate its generalization, we evaluate both Qwen2.5-VL-72B and Gemini 2.0 Flash, showing that MAPS performs well across models of varying scales and capabilities.
+Figure 6 presents results for three sets of base models.
+In each set, we compare MLLMs and MAPS on mathematical, physical, and chemical sub-tasks.
+MAPS consistently outperforms the base models.
+For example, MAPSQwen improves Qwen2.5-VL-72B by 12.4% in physics, while MAPSGemini improves Gemini by 4.2%.
+Similar gains are observed in math and chemistry, demonstrating MAPS’s effectiveness on both open-source and closed-source MLLMs.
+Time Efficiency
+Simpler formats and lower difficulty yield faster solving times.
+Solving time efficiency varies by question type, answer type, category, and difficulty, with multiple-choice and integer-type questions being the fastest, while higher difficulties and complex formats require more time.
+Figure 7 illustrates the solving time efficiency across various dimensions—question types, answer formats, subject categories, and difficulty levels—with all times normalized to a 100s benchmark.
+Predefined structure and conceptual simplicity reduce reasoning time.
+Multiple-choice questions are solved more quickly thanks to predefined answer options that limit the need for extensive reasoning or exploration.
+Integer-type answers also show high efficiency, often tied to simpler arithmetic or structured formats requiring minimal inference.
+General category questions are faster on average, likely due to lower conceptual and reasoning complexity compared to domain-specific tasks.
+In contrast, open-ended questions demand deeper analysis and justification, leading to longer solving times.
+Finally, solving efficiency declines with increased difficulty: as question complexity rises, so does the required reasoning time, reflecting greater cognitive and computational demands.
 Figure 6: Performance Comparison of MAPS on Math, Physics, and Chemistry Subtasks in the EMMA Dataset with GPT-4o, Gemini, and Qwen2.5-VL-72B as Bases.
-The removal of the Critic agent causes the smallest performance loss.
-It results in only a 7.05% decrease, underscoring its role in providing feedback and corrections.
-While this mechanism allows MAPS to backtrack and refine its reasoning, its impact is less significant than that of other agents.
-Removing the Scholar agent results in 11.49% performance drops, highlighting the importance of searching and integrating domain-specific knowledge.
-Finally, the removal of the Aligner agent causes a 10.86% drop, indicating that while diagram and context alignment is valuable, its effect is smaller compared to other components.
-4.2 Base Model Generalization
-MAPS can improve performance across different base models.
-We conduct a series of experiments to verify whether our MAPS framework can demonstrate robust generalization across different base LLMs.
-The results showcase the model’s robustness and transferability, further affirming the high adaptability of the MAPS framework, which achieves consistently excellent performance across various foundational models.
-In particular, this underscores the universality and flexibility of MAPS.
-To further validate its generalization across different models, we specifically evaluate another base model, Qwen2.5-VL-72B, as well as Gemini 2.0 Flash, thereby demonstrating that MAPS performs outstandingly across models of varying scales and capabilities.
-Figure 5 presents the experimental results of three sets of base models.
-In each set, the performance of mathematical, physical, and chemical sub-tasks is compared between MLLMs and MAPS with the respective base model.
-The results show varying degrees of improvement for each task.
-For example, in the physical task, $MAPS_{Qwen}$, based on Qwen2.5-VL-72B, shows a 12.4% improvement compared to Qwen2.5-VL-72B, while $MAPS_{Gemini}$, based on Gemini, shows a 4.2% improvement compared to Gemini.
-Similar improvements are observed in the mathematical and chemical sub-tasks as well.
-This indicates that MAPS can extend effectively on both closed-source and open-source MLLMs and achieve good experimental results across different base models of MLLMs.
 Figure 7: An analysis of the solving time efficiency across different question types, answer types, question categories, and question difficulties.
-4.3 Time Efficiency
-Solving time efficiency varies by question type, answer type, category, and difficulty, with multiple-choice and integer-type questions being the fastest, while higher difficulties and more complex questions require more time.
-Figure 6 presents the solving time efficiency under different question types, answer types, categories, and difficulties, with all times represented relative to a 100s benchmark.
-From the graphs, it is clear that multiple-choice questions are solved faster as the options are already provided, making the process more time-efficient.
-In contrast, open-ended questions take longer due to the lack of predefined answers and the need for deeper reasoning and analysis.
-Integer-type answers show the highest efficiency, as they often correspond to simpler questions or multiple-choice tasks that involve less reasoning.
-General questions also have the shortest solving times, as they typically require less complex reasoning compared to more specialized problem types.
-Finally, higher difficult levels lead to a decrease in solving time efficiency, as more complex problems require more time for deeper reasoning and analysis.
-5 Related Works
-The related work is structured into two main aspects: first, an introduction to MSPs solving; and second, an exploration of multi-agent techniques.
-MSPs Solving.
-The research of MSPs spans across multiple fields, including mathematics, physics, and chemistry, with each area focusing on enhancing problem-solving abilities.
-In mathematics, studies [9, 11, 34] explore various methods to improve mathematical problem-solving, such as algorithm optimization, educational strategies, and the use of artificial intelligence.
-These approaches aim to boost the efficiency, accuracy, and depth of mathematical reasoning, fostering innovation and progress in the field.
-In the field of physics, the papers [24, 17, 3] emphasize the integration of different information types, such as images and text, through multimodal learning to enhance the efficiency and precision of problem solving.
-By utilizing techniques like reinforcement learning, generative models, and intelligent prompting, these studies optimize models’ performance in tackling complex physics problems.
-In chemistry, three articles [1, 18, 21] investigate the role of multimodal learning in solving chemical problems.
-By combining diverse information sources, including images and text, and employing techniques such as generative models and molecular geometry reasoning, they aim to improve both the efficiency and accuracy of solving chemistry problems, driving innovation in chemistry education and research.
+Related Works
+The related work is structured into two main aspects: Firstly, an introduction to complex problem reasoning; Secondly, an exploration of multi-agent techniques.
+Complex Problem Reasoning.
+The research of complex problem reasoning spans across multiple fields, including mathematics, physics, and chemistry, with each area focusing on enhancing problem-solving abilities.
+In mathematics, studies (didolkar2024metacognitive; fitriana2024influence; tong2025dart) explore various methods to improve mathematical problem-solving, such as algorithm optimization, educational strategies, and the use of artificial intelligence.
+These approaches aim to boost the efficiency, accuracy, and depth of mathematical reasoning.
+In the field of physics, the papers (mustofa2024utilizing; kapuriya2024mm; anand2024mm; wu2024gps) emphasize the integration of different information types, such as images and text, through multimodal learning to enhance the efficiency and precision of problem solving.
+In chemistry, three articles (alasadi2024multimodal; kiernan2024resources; li2024chemvlm; lin2025foundation; dang2025disentangled) investigate the role of multimodal learning in solving chemical problems.
+By combining diverse information sources, including images and text, and employing techniques such as generative models and molecular geometry reasoning, they aim to improve both the efficiency and accuracy of solving chemistry problems.
 Multi-Agent.
-Multi-agent systems, built on LLMs, consist of multiple AI agents that specialize in specific tasks, working together to solve complex problems [30, 42, 39, 32].
+Multi-agent systems, built on LLMs, consist of multiple AI agents that specialize in specific tasks, working together to solve complex problems (richards2023auto; yang2023auto; wu2023autogen; sun2023corex; zhang2025maps; zhang2025mars; zhang2025gkg; yan2025mur).
 When presented with a problem, these agents decompose it into smaller, manageable subtasks and utilize various tools, such as internet data retrieval, to solve them through iterative steps.
-Several studies [28, 37, 40, 26] have employed multi-agent systems to tackle challenges like problem identification, code writing and debugging, data visualization, and providing interactive feedback to human users.
-In their work, Ni and Buehler [25] highlights the potential of AI-driven multi-agent teams in solving mechanical problems autonomously, demonstrating an enhanced capability for understanding, formulating, and validating engineering solutions through self-correction and collaborative refinement.
-Inspired by the research, we developed the MAPS method, which leverages multi-agent collaborative learning and stepwise problem-solving to provide innovative solutions for MSPs.
+Several studies (poldrack2023ai; wang2024survey; xi2025rise; ni2021deep) have employed multi-agent systems to tackle challenges like problem identification, code writing and debugging, data visualization, and providing interactive feedback to human users.
+In their work, ni2024mechagents highlights the potential of AI-driven multi-agent teams in solving mechanical problems autonomously, demonstrating an enhanced capability for understanding, formulating, and validating engineering solutions through self-correction and collaborative refinement.
+Inspired by the research, we developed the MAPS method, which leverages multi-agent collaborative learning and stepwise problem-solving to provide innovative solutions for complex problem reasoning.
 By combining the strengths of AI agents, complex problems can be broken down into subtasks and solved step by step through collaboration, improving efficiency and accuracy.
-6 Conclusion
-This study introduces a MAPS approach utilizing a multi-agent framework based on the Big Seven Personality theory and Socratic guidance to tackle the challenges of multi-modal comprehensive reasoning and the lack of reflective capabilities.
-The framework involves seven agents, each specializing in distinct aspects of problem-solving.
+Conclusion
+This study presents MAPS, a multi-agent framework grounded in the Big Five Personality Theory and guided by Socratic principles, designed to address the challenges of multimodal comprehensive reasoning and enhance reflective capabilities.
+The framework involves five agents, each specializing in distinct aspects of problem-solving.
 To address the first challenge, a four-agent strategy is proposed, where each agent focuses on specific stages of the reasoning process.
 Additionally, the Critic agent addresses the second challenge through Socratic reflection and critical feedback.
-Extensive experiments on the EMMA, Olympiad, and MathVista datasets validate MAPS’s effectiveness in overcoming these issues and enhancing performance across various reasoning tasks.
+Extensive experiments on the EMMA, OlympiadBench, and MathVista datasets validate MAPS’s effectiveness in overcoming these issues and enhancing performance across various reasoning tasks.
 Meanwhile, we perform additional analytical experiments to assess the model’s advancement as well as its generalization.
-Supplementary Material
-Appendix A Datasets and Tasks
-This study utilizes the latest three MSPs datasets, namely MathVista [23], OlympiadBench [16], and EMMA [14].
-Tasks ABBR. Test
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (No.
+62137002, 62277042, 62293553, 62450005, 62437002, 62477036, 62477037, 62176209, 62192781, 62306229), the “LENOVO-XJTU” Intelligent Industry Joint Laboratory Project, the Shaanxi Provincial Social Science Foundation Project (No.
+2024P041), the Natural Science Basic Research Program of Shaanxi (No.
+2023-JC-YB-593), and the Youth Innovation Team of Shaanxi Universities “Multi-modal Data Mining and Fusion”.
+Appendix A Proof of Proposition
+Proof of Proposition 1 (Monotonic Free-Energy Descent)
+Statement recap. Let $F^{(t)}$ denote the variational free energy after the $t$-th MAPS iteration:
+$F^{(t)}=\mathbb{E}_{q^{(t)}(\theta)}[-\log p(\mathbf{x},a_{i}\mid\theta)]+\mathrm{KL}(q^{(t)}(\theta)\|p(\theta)),$ (14)
+where $q^{(t)}(\theta)$ is the approximate posterior induced by the agent outputs at iteration $t$. The proposition claims that each Critic-guided update ensures
+$F^{(t+1)}\leq F^{(t)},$ (15)
+which implies that $\{F^{(t)}\}_{t=0}^{\infty}$ is a non-increasing sequence that converges.
+Step 1: Free Energy as a Variational Upper Bound. By standard variational inference, the marginal log-likelihood satisfies
+$\log p(\mathbf{x},a_{i})=\mathcal{L}(q)+\mathrm{KL}(q(\theta)\|p(\theta\mid\mathbf{x},a_{i})),$ (16)
+where the ELBO is defined as
+$\mathcal{L}(q)=\mathbb{E}_{q(\theta)}[\log p(\mathbf{x},a_{i}\mid\theta)]-\mathrm{KL}(q(\theta)\|p(\theta)).$ (17)
+This leads to the variational free energy $F(q):=-\mathcal{L}(q)$, which upper-bounds the negative log marginal likelihood:
+$F(q)\geq-\log p(\mathbf{x},a_{i}).$ (18)
+Step 2: Critic-Guided Local Update.
+Suppose the Critic identifies the weakest stage $k^{*}=\arg\min_{k}f_{k}^{(t)}$ and re-executes $\mathcal{A}_{k^{*}}$, yielding updated output $\mathcal{S}_{k^{*}}^{(t+1)}$.
+The updated trajectory leads to a refined posterior:
+$q^{(t+1)}(\theta)=\mathcal{G}(q^{(t)}(\theta),\mathcal{S}_{k^{*}}^{(t+1)}),$ (19)
+where $\mathcal{G}$ represents the update mechanism that propagates changes forward through dependent modules.
+As the re-computed stage corrects suboptimal reasoning or improves knowledge alignment, it leads to a strictly better or equal ELBO, and thus
+$F^{(t+1)}=F(q^{(t+1)})\leq F(q^{(t)})=F^{(t)}.$ (20)
+Step 3: Convergence. Since $F^{(t)}$ is lower-bounded and non-increasing, the sequence converges:
+$\lim_{t\to\infty}F^{(t)}=F^{(\infty)}.$ (21)
+Proof of Proposition 3 (Collaborative Information Bottleneck)
+Statement recap.
+Given input $\mathbf{x}=(d_{i},c_{i},q_{i})$ and target answer $a_{i}$, let $\mathcal{S}_{k}$ denote the intermediate output of the $k$-th agent in the MAPS reasoning process.
+Then the objective is to find agents that minimize cumulative information compression from input while preserving task-relevant content:
+$\min_{\{\mathcal{A}_{k}\}}\sum_{k=1}^{4}I(\mathbf{x};\mathcal{S}_{k})\quad\text{s.t.}\quad I(\mathcal{S}_{k};a_{i})\geq\varepsilon,$ (22)
+where $I(\cdot;\cdot)$ denotes mutual information, and $\varepsilon>0$ is a threshold for information sufficiency.
+Step 1: Motivation from the Information Bottleneck (IB) Principle.
+In standard IB theory (tishby2000information), one seeks a compressed representation $\mathcal{S}$ of input $\mathbf{x}$ that retains information about target $a_{i}$.
+The trade-off objective is:
+$\min I(\mathbf{x};\mathcal{S})\quad\text{subject to}\quad I(\mathcal{S};a_{i})\geq\varepsilon.$ (23)
+MAPS generalizes this to a multi-agent, multi-stage setting, where each agent $\mathcal{A}_{k}$ produces $\mathcal{S}_{k}$, and the full reasoning trajectory is $\mathcal{T}=\{\mathcal{S}_{1},\dots,\mathcal{S}_{4}\}$.
+The objective thus becomes collaborative:
+$\min\sum_{k=1}^{4}I(\mathbf{x};\mathcal{S}_{k})\quad\text{s.t.}\quad\forall k,\;I(\mathcal{S}_{k};a_{i})\geq\varepsilon.$ (24)
+This reflects two goals: (1) compress redundant input features per stage, and (2) preserve task-relevant evidence at each step.
+Step 2: Role of the Critic as a Constraint Monitor.
+The Critic observes the reasoning trajectory $\mathcal{T}$ and computes scores $\mathbf{f}=\{f_{1},f_{2},f_{3},f_{4}\}$, where each $f_{k}$ implicitly estimates $I(\mathcal{S}_{k};a_{i})$.
+If $f_{k}<\tau$, we assume the sufficiency constraint is violated:
+[…]
+The revision process updates $\mathcal{S}_{k}$ to $\mathcal{S}_{k}^{\prime}$, ideally increasing $I(\mathcal{S}_{k}^{\prime};a_{i})$ while not increasing $I(\mathbf{x};\mathcal{S}_{k}^{\prime})$.
+Since the Critic only intervenes when constraints are violated, each update maintains feasibility and possibly reduces the objective.
+Step 3: Convergence and Validity.
+As each update either preserves or improves the sufficiency $I(\mathcal{S}_{k};a_{i})$, and the overall input compression is bounded below by zero, the process converges to a locally optimal set of agents $\{\mathcal{A}_{k}^{*}\}$ satisfying:
+$\forall k,\quad I(\mathcal{S}_{k}^{*};a_{i})\geq\varepsilon,\quad\text{with minimal}\quad\sum_{k}I(\mathbf{x};\mathcal{S}_{k}^{*}).$ (26)
+This completes the proof that the MAPS four-step process can be viewed as a constrained collaborative information bottleneck optimization.
+Appendix B Datasets
+This study utilizes the latest three multimodal reasoning datasets, namely MathVista, OlympiadBench, and EMMA.
+Dataset ABBR. Size
 MathVista
@@
 EMMA
-       Mathmatics Math 100
+     Math 100
      Physics Phy. 100
@@
 The ‘ABBR.’ column represents the abbreviations for all the tasks.
+MathVista.
 MathVista is a large-scale scientific reasoning dataset that spans two subdomains: mathematics and general, aiming to assess the comprehensive capabilities of machine learning models in solving complex scientific problems.
@@
 The release of MathVista supports interdisciplinary scientific research.
+OlympiadBench.
 OlympiadBench consists of two subdomains, mathematics and physics, and is specifically designed for Mathematical and Physical Olympiads, featuring a wide range of challenging problems to assess models’ performance on high-level scientific tasks.
@@
 To ensure data balance, 300 samples were taken from both the Chinese college-level mathematics and physics subsets.
+EMMA.
 EMMA is a multimodal scientific reasoning dataset covering three subdomains: mathematics, physics, and chemistry.
@@
 This version uses the EMMA dataset, which contains 100 data points from each subdomain (mathematics, physics, and chemistry).
-Appendix B Experiment Settings and Baselines
-We select GPT-4o [^3] https://openai.com/api/, a powerful MLLM, as our primary agent for solving MSPs tasks.
+Appendix C Experiment Settings and Baselines
+We select GPT-4o (achiam2023gpt), a powerful MLLM, as our primary agent for complex problem reasoning.
 GPT-4o not only demonstrates strong reasoning and generation capabilities across a wide range of multimodal processing tasks, but also excels in efficiently exploring multiple perspectives when faced with complex scientific domain requirements.
@@
 We use accuracy as our primary evaluation metric to comprehensively assess the performance of different methods across diverse task scenarios.
-The experimental results present in Table 1 offer a thorough comparison of the performance of MAPS against all baseline methods.
-The experiments are primarily conducted on three MSPs datasets in  Appendix A, where we provide a detailed comparison of MAPS and the baseline models across four types of tasks: mathematics, physics, chemistry, and general tasks.
+The experimental results offer a thorough comparison of the performance of MAPS against all baseline methods.
+The experiments are primarily conducted on three datasets in Appendix B, where we provide a detailed comparison of MAPS and the baseline models across four types of tasks: mathematics, physics, chemistry, and general tasks.
 Our approach achieves a new SOTA performance.
-To further strengthen the comparison,  Appendix C includes a generalization experiment conducted on the physics data subset of the DiagramQG [^4] https://huggingface.co/datasets/zhibei1204/DiagramQG, which further demonstrates the robustness and effectiveness of our model.
-Figure 7: The generalization experiments conducted on the DiagramQG physical dataset, which are based on the GPT-4o base model and the incremental part of MAPS.
+To further strengthen the comparison, Appendix D includes a generalization experiment conducted on the physics data subset of the DiagramQG (zhang2025diagram), which further demonstrates the robustness and effectiveness of our model.
 We compare MAPS with three categories of baseline methods: original baselines, direct approach-based strong baselines, and CoT-enhanced strong baselines.
@@
 (2) The direct approach-based strong baselines include some of the most powerful closed-source and open-source large language models (MLLMs) currently available worldwide.
-These include Claude 3.5 Sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini 2.0 Flash [^6] https://deepmind.google/technologies/gemini/flash/, GPT-4o, Qwen2.5-VL-72B [^7] https://huggingface.co/Qwen/Qwen2.5-VL-72B-Instruct, InternVL2.5-8B-MPO [^8] https://huggingface.co/OpenGVLab/InternVL2-8B-MPO, and LLaVA-Onevision-72B [^9] https://huggingface.co/llava-hf/llava-onevision-qwen2-72b-ov-hf.
-(3) To ensure a fair comparison, the third category of baselines builds on the second by adding Chain-of-Thought (CoT) reasoning, which aims to enhance the capabilities of the strong MLLMs from the second category [38].
-Appendix C Generalization Experiments
+These include Claude 3.5 Sonnet (kevian2024capabilities), Gemini 2.0 Flash (comanici2025gemini), GPT-4o, Qwen2.5-VL-72B (bai2023qwen), InternVL2.5-8B-MPO (chen2024expanding), and LLaVA-Onevision-72B (li2024llava).
+(3) To ensure a fair comparison, the third category of baselines builds on the second by adding Chain-of-Thought (CoT) reasoning, which aims to enhance the capabilities of the strong MLLMs from the second category.
+Appendix D Generalization Experiments
+Figure 8: The generalization experiments conducted on the DiagramQG physical dataset, which are based on the GPT-4o base model and the incremental part of MAPS.
+Figure 9: A case study of a specific solving process, illustrating the detailed steps involved in solving the problem.
+This includes the various stages of problem-solving as well as the feedback and backtracking mechanisms that help refine and improve the solution.
 To further validate the generalization ability of MAPS, we conducted detailed experiments on the physical subset of the DiagramQG dataset.
@@
 These experiments clearly highlight the strong adaptability of MAPS across different subsets of the dataset.
-Across various question categories in the DiagramQG physical dataset, MAPS achieved a significant performance boost, with the maximum improvement reaching 19.51% and an overall improvement of 7.71%.
-These results not only demonstrate the superiority of MAPS but also indicate that its Seven Personality Agents architecture has a strong generalization ability, enabling it to maintain excellent performance across different datasets and tasks.
+Across various question categories in the DiagramQG physical dataset, MAPS achieved a performance boost, with the maximum improvement reaching 19.51% and an overall improvement of 7.71%.
+These results not only demonstrate the superiority of MAPS but also indicate that its Five Personality Agents architecture has a strong generalization ability, enabling it to maintain excellent performance across different datasets and tasks.
 This provides strong support for the practical application of MAPS, showcasing its potential in tackling complex tasks.
-Figure 8: A case study of a specific solving process, illustrating the detailed steps involved in solving the problem.
-This includes the various stages of problem-solving as well as the feedback and backtracking mechanisms that help refine and improve the solution.
-Appendix D Full-Process of MAPS
+Appendix E Full Process of MAPS
 Figure 9 illustrates the four-step solving process along with the feedback process from the Interpreter, Aligner, Scholar, and Critic agents, using a multimodal physics problem from the EMMA dataset.
@@
 Through the feedback mechanism of the Critic agent, MAPS is able to identify and correct potential errors or shortcomings at each step, thus enhancing the overall accuracy and reliability of the solution process.
-Figure 9: A complete example of the collaborative output from all agents in an iteration, using the multimodal physics problem from the OlympiadBench dataset.
-This example demonstrates how each agent contributes to the problem-solving process, collaborating to produce a refined solution step by step.
-Figure 9 presents a comprehensive step-by-step process for solving a MSPs task, using a multimodal physics problem from the EMMA dataset.
-As illustrated in  Figure 3, this example demonstrates the entire process of solving a physics problem.
-By utilizing MAPS agents based on the Big Seven Personality Model, the system engages in collaborative learning, progressively solving the problem and arriving at the correct final answer.
-From the Manager, which specifically plans the entire MSPs solving process and feedback strategies, to the UserProxy that accepts inputs and task descriptions, the solution process is carried out in four key steps.
+Figure 10: A complete example of collaborative output from all agents in an iteration, using a multimodal physics problem from OlympiadBench involving the axial electric field of a dipole in the far-field.
+This illustrates how agents work together step by step to refine the solution.
+Figure 10 presents a comprehensive step-by-step process for complex problem reasoning, using a multimodal physics problem from the EMMA dataset.
+This example demonstrates the entire process of solving a physics problem.
+By utilizing MAPS agents based on the Big Five Personality Model, the system engages in collaborative learning, progressively solving the problem and arriving at the correct final answer.
+The solution process is carried out in four key steps.
 These include the Interpreter, Aligner, Scholar, and Solver, each responsible for gradually refining the solution step-by-step.
@@
 Figure 10 sequentially illustrates the roles of each agent and their collaboration, showcasing how problem solving and optimization are effectively executed at every stage, ensuring both the accuracy and efficiency of the final result.
-Appendix E Prompts for Agents
-Tables 4, 5 and 6 provides a summary of the prompts used for each agent in this paper, with each agent playing a pivotal role in the overall problem-solving process.
-The process begins with the Manager, which is responsible for strategically planning the entire MSPs solving procedure and feedback mechanisms.
-The Manager defines the high-level flow, ensuring that all agents operate cohesively.
-The UserProxy agent then takes over by accepting the user inputs and task descriptions, forming the basis for the entire problem-solving process.
-Following this, the solution is gradually refined by four core agents: Interpreter, Aligner, Scholar, and Solver.
+Appendix F Prompts for Agents
+Table 45, and 6 provides a summary of the prompts used for each agent in this paper, with each agent playing a pivotal role in the overall problem-solving process.
+The solution is gradually refined by four core agents: Interpreter, Aligner, Scholar, and Solver.
 The Interpreter first processes the task description, breaking down and understanding the problem.
@@
 The Critic then suggests modifications and backtracks to revise the solution, ensuring the process remains optimized and the final result is both accurate and robust.
-Manager
-You are a task manager, responsible for managing and deciding the order of each Agent’s output, ensuring that tasks are executed in the correct sequence.
-You will determine when each step should be executed based on the task requirements and coordinate the outputs of each Agent.
-Your goal is to ensure that the task process is efficient and orderly, and adjust the execution priority of steps when necessary.
-UserProxy
-Here are some questions that require careful thought, so please think deeply, solve the questions carefully, and output your answers.
 Interpreter
@@
 [Relevance Filtering] ”Through lexical-semantic mapping, which aligned components directly correspond to the question’s: 1) Key inquiry points 2) Required evidence types 3) Implicit knowledge domains 4) Potential inference pathways?”
-Table 4: A summary of the prompts used by the Manager, UserProxy, Interpreter, and Aligner agents in this paper.
+Table 4: A summary of the prompts used by the Interpreter, and Aligner agents in this paper.
 scholar